### v1

# HARec: Hyperbolic Graph-LLM Alignment for Exploration and Exploitation in Recommender Systems

Abstract.
Modern recommendation systems often create information cocoons, limiting users’ exposure to diverse content. To enhance user experience, a crucial challenge is developing systems that can balance content exploration and exploitation, allowing users to adjust their recommendation preferences. Intuitively, this balance can be achieved through a tree-structured representation, where depth search facilitates exploitation and breadth search enables exploration. However, current works face two challenges to achieve this target: (1) Euclidean methods fail to fully capture hierarchical structures and lack flexibility in balancing exploration-exploitation, while (2) hyperbolic approaches, despite better hierarchical modeling, suffer from insufficient semantic alignment due to their reliance on Euclidean text encoders. To address these challenges, we propose HARec, a hyperbolic representation learning framework that jointly aligns user-item collaborative information with textual descriptions in hyperbolic space. Our framework introduces two key technique novelty: (1) a hierarchical-aware graph-llm alignment mechanism that enables better hierarchical representation, and (2) a hyperbolic hierarchical tree structure that facilitates user-adjustable exploration-exploitation trade-offs. Extensive experiments demonstrate that HARec consistently outperforms both Euclidean and hyperbolic baselines, achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics.
Recommender Systems, Representation Learning, Hyperbolic Space
Authors affiliated with Snap Inc. served in advisory roles only for this work.
Corresponding author
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; Aug 3–Aug 7, 2025; Woodstock, NYccs: Information systems Recommender systemsccs: Information systems Collaborative filteringccs: Mathematics of computing Geometric topologyccs: Information systems Personalization
1. INTRODUCTION
As online platforms grow exponentially, users are overwhelmed by an influx of information, resulting in information cocoons that restrict their exposure to diverse and novel content (Kunaver and Požrl, 2017; Castells et al., 2021).
Traditional recommendation approaches have primarily relied on collaborative filtering and interaction data (He et al., 2020; Koren et al., 2009), where users with similar historical preferences are expected to share future interests (Chen et al., 2021). However, these methods face two critical challenges in improving user experience. First, they often fail to capture the rich semantic relationships between items and evolving user preferences. For example, a jazz enthusiast might wish to explore rock music, yet current models struggle to understand these semantic connections and adapt to such evolving preferences (as illustrated in Figure 1). Second, existing approaches inadequately capture the underlying hierarchical structure inherent in user-item networks (Zheng et al., 2017; Unger and Tuzhilin, 2020), a critical factor for improving performance and personalization.
The emergence of large language models (LLMs) offers a solution to the limitations of semantic understanding. However, integrating LLM-derived semantic features with traditional collaborative signals remains challenging, especially when capturing the hierarchical structure of user preferences and item relationships.
Hyperbolic space has emerged as a promising solution to these issues. It is a powerful tool for modeling tree-like structures and power-law distributions (Krioukov et al., 2010; Sun et al., 2021; Yang et al., 2022b, a), where its exponential volume growth naturally aligns with hierarchical, scale-free structures (Ganea et al., 2018; Chami et al., 2019; Liu et al., 2019; Zhang et al., 2021; Yang et al., 2023). In hyperbolic space, distances increase exponentially from the origin, allowing popular items or exploratory users to be positioned closer to the origin, while niche items or users with focused interests spread toward the boundaries. This spatial configuration enables hyperbolic graph neural network (GNN) models to perceive hierarchical relationships effectively, enhancing recommendation performance (Sun et al., 2021; Yang et al., 2022a, b).
However, current hyperbolic recommendation systems still face two key limitations. First, hyperbolic approaches often lack proper semantic alignment due to their dependence on Euclidean text encoders, which struggle to integrate within hyperbolic space. This mismatch leaves existing hyperbolic models vulnerable to noise in collaborative information, as they fail to capture semantic insights effectively, impacting their robustness and accuracy. Second, hyperbolic recommender systems lack a systematic approach to balancing exploration (introducing new content) and exploitation (focusing on existing preferences) (Wilson et al., 2021). This limitation hinders their adaptability to users’ evolving interests and reduces the diversity and relevance of recommended items.
Proposed Methods. To address these challenges, we propose HARec, a hyperbolic representation learning framework that facilitates exploration-exploitation through an intuitive tree-structured representation, where depth search enables exploitation and breadth search allows exploration. Our framework introduces two key components: (1) a hierarchical-aware graph-llm alignment mechanism that enhances semantic understanding and reduces noise sensitivity by jointly aligning textual descriptions with user-item collaborative information in hyperbolic space, and (2) a hyperbolic hierarchical tree structure that enables user-adjustable exploration-exploitation trade-offs. This tree structure organizes preferences hierarchically, with higher layers representing broader interests, allowing systematic navigation between exploration and exploitation.
To validate the effectiveness of HARec, we conducted extensive experiments across multiple dimensions. Our model consistently outperformed baselines in both utility and diversity, marking a significant achievement as it is the first model to excel in both aspects simultaneously. To further highlight the benefits of our hyperbolic alignment framework, we compared HARec with feature-empowered baselines, where it demonstrated superior performance. Additionally, we conducted detailed analysis on head and tail items, showing that HARec effectively boosts recommendations for tail items, thus enhancing diversity. An ablation study, along with a decoder module analysis, confirms HARec’s capability to leverage both semantic meaning and collaborative information, providing deeper insights into the advantages of our approach.
Contributions. To summarize, our primary contribution are in three aspects. The first is the proposed HARec, a hyperbolic representation learning framework that effectively balances exploration and exploitation by integrating semantic and collaborative information, supported by theoretical analysis demonstrating the advantages of this integration in hyperbolic space. In addition, we design a novel mechanism enabling personalized control over the exploration-exploitation trade-off through a hierarchical tree structure in hyperbolic space, allowing users to dynamically adjust their recommendation preferences. Moreover, through extensive experiments evaluating both accuracy and diversity metrics, we demonstrate that HARec achieves state-of-the-art performance across both criteria - achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics, making it the first model to excel in both criteria simultaneously.
Figure 1. Exploration and exploitation in hyperbolic space.
2. Related Work
In this section, we review commonly used graph-based collaborative filtering approaches, large language models (LLMs) for recommendation, and hyperbolic representation learning methods.
Graph Collaborative Filtering. Graph Collaborative Filtering has emerged as a dominant approach in recommender systems, with graph neural networks (GNNs) playing a central role. Unlike traditional collaborative filtering techniques, such as matrix factorization (Koren et al., 2009), GNNs capitalize on the relational structure inherent in user-item interactions, enabling a more nuanced understanding of complex dependencies. Prominent graph-based methods, such as NGCF (Wang et al., 2019), SGL (Wu et al., 2021), and LightGCN (He et al., 2020), have been widely adopted and have consistently achieved state-of-the-art performance, significantly boosting recommendation accuracy. Nevertheless, most of these models are constrained to Euclidean space, which may fall short in capturing the underlying power-law distribution commonly seen in user-item networks.
Large Language Models for Recommendation. Large language models (LLMs) are increasingly used in recommender systems due to their strong capability in understanding textual information. Some approaches leverage LLMs directly for various downstream recommendation tasks (Bao et al., 2023; Li et al., 2023; Lyu et al., 2023; Zheng et al., 2024; Zhang et al., 2023), while others utilize LLMs as an auxiliary tool for collaborative filtering (Ren et al., 2024; Xi et al., 2024; Wei et al., 2024). However, both approaches often require substantial inference time and are computationally inefficient. In our HARec, we use language models solely to enhance data quality, avoiding the need for repeated inference during each recommendation by relying on hyperbolic embeddings for the final output.
Hyperbolic Representation Learning. To address the limitations of Euclidean space in modeling hierarchical structures, researchers have investigated the use of hyperbolic space (Chami et al., 2019; Liu et al., 2019; Yang et al., 2022b; Zhang et al., 2021), where distances grow exponentially as embeddings move away from the origin. This property makes hyperbolic geometry particularly well-suited for modeling user-item interactions by more accurately capturing latent hierarchies and dependencies. Consequently, hyperbolic representation learning has gained considerable attention in the recommender systems domain (Ying et al., 2018; Vinh Tran et al., 2020; Sun et al., 2021; Yang et al., 2022a), demonstrating improved performance over traditional Euclidean models.
However, none of these hyperbolic models fully leverage the rich semantic information embedded in language-based datasets. While some approaches incorporate knowledge graphs as supplementary information (Tai et al., 2021; Du et al., 2022), these models primarily focus on refining hierarchical structures rather than achieving a deeper semantic understanding of each item or user.
3. PRELIMINARIES
Embeddings learned from the above graph collaborative filtering are by default in the Euclidean space, which is unable to capture tree-like structures. To address the problem, we introduce how to derive these embeddings in a hyperbolic space.
Riemannian Manifold. A Riemannian manifold $(\mathcal{M},g)$ is a smooth manifold that features an inner product $g_{x}:T_{x}\mathcal{M}\times T_{x}\mathcal{M}\rightarrow\mathbb{R}$ at each point $x\in\mathcal{M}$. This inner product, known as the Riemannian metric, defines geometric characteristics of the space such as angles and curve lengths. The curvature of Riemannian manifolds results in different types of geometries: elliptic geometry with positive curvature, Euclidean geometry with zero curvature, and hyperbolic geometry with negative curvature. In this work, we focus on hyperbolic geometry, characterized by negative curvature. Hyperbolic space can be described by several equivalent models, each offering unique properties while being mathematically equivalent. The Lorentz model, also referred to as the hyperboloid model, is one of the commonly used representations of hyperbolic space.
Lorentz Model. An $n$-dimensional Lorentz manifold with negative curvature $-1/\kappa$ $(\kappa>0)$ is described as the Riemannian manifold $(\mathbb{H}^{n}_{\kappa},g_{\mathcal{L}})$. Here, $\mathbb{H}^{n}_{\kappa}=\{x\in\mathbb{R}^{n+1}:\langle x,x\rangle_{\mathcal{L}}=-\kappa,x_{0}>0\}$, $g_{\mathcal{L}}=\eta$ with $\eta=\mathbf{I}_{n}$ except $\eta_{0,0}=-1$, and $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ represents the Lorentzian inner product, which is defined as by
(1) $\langle x,y\rangle_{\mathcal{L}}:=-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}.$
A tangent space is an $n$-dimensional vector space that approximates Euclidean space to hyperbolic space. Specifically, given a point $x\in\mathbb{H}_{\kappa}^{n}$, there exists a tangent space $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$ that approximates $\mathbb{H}_{\kappa}^{n}$:
(2) $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}:=\{v\in\mathbb{R}^{n+1}:\langle v,x\rangle_{\mathcal{L}}=0\}.$
The tangent space is defined to be orthogonal to $x$ with respect to the Lorentzian inner product, therefore making the approximation accurate at the reference point.
We also need to define transitions between $\mathbb{H}_{\kappa}^{n}$ and $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$, which give rise to the exponential and logarithmic maps. For $x\in\mathbb{H}_{\kappa}^{n}$ and $v\in\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$ with $v\neq\mathbf{0}$ and $y\neq x$, there exists a unique geodesic $\gamma:[0,1]\to\mathbb{H}_{\kappa}^{n}$ such that $\gamma(0)=x$ and $\gamma^{\prime}(0)=v$. The exponential map $\exp_{x}:\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}\to\mathbb{H}_{\kappa}^{n}$ is given by $\exp_{x}(v)=\gamma(1)$:
(3) $\exp_{x}^{\kappa}(v)=\cosh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)x+\sqrt{\kappa}\sinh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)\frac{v}{\|v\|_{\mathcal{L}}},$
where $\|v\|_{\mathcal{L}}=\sqrt{\langle v,v\rangle_{\mathcal{L}}}$ denotes the Lorentzian norm of $v$. The logarithmic map $\log_{x}$ is the inverse of the exponential map $\exp_{x}$:
(4) $\log_{x}^{\kappa}(y)=d_{\mathcal{L}}^{\kappa}(x,y)\frac{y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x}{\|y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x\|},$
where $d_{\mathcal{H}}^{\kappa}(x,y)$ represents the distance between $x$ and $y$ in $\mathbb{H}^{n}_{\kappa}$:
(5) $d_{\mathcal{H}}^{\kappa}(x,y)=\sqrt{\kappa}\,\text{arcosh}\left(-\langle x,y\rangle_{\mathcal{L}}/\kappa\right).$
We choose $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ as the reference point for all the operations, making the transitions reproducible. In this work, we set $\kappa=1$, resulting in a curvature of $-1$. Therefore, in the following discussions, we would omit $\kappa$ for simplicity.
4. METHODOLOGY
In this section, we present a comprehensive overview of our proposed model, HARec. Our approach not only aims to deliver accurate recommendations but also addresses users’ growing demand for more diverse suggestions. Furthermore, by constructing a hierarchical tree from the learned embeddings, we introduce an additional mechanism that enables users to control the degree of exploration in the recommended items. The overall framework is in Figure 2.
Figure 2. The overall architecture of HARec. (i) Hyperbolic Graph Collaborative Filtering: Encodes collaborative information using hyperbolic GNNs; (ii) Hyperbolic Alignment: Aligns semantic and collaborative information within hyperbolic space; (iii) Hyperbolic Hierarchical Structure: Builds hierarchy tree from hyperbolic embeddings.
4.1. Hyperbolic Graph Collaborative Filtering
Similar to Euclidean space, hyperbolic graph collaborative filtering captures high-order dependencies between users and items via a message aggregation mechanism. In our method, we use the Lorentz representation for both users and items.
Hyperbolic Embedding Initialization. As discussed in section 3, we fix the origin $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ and use it as the reference point. We adopt the Gaussian sampling method for the initialization of embeddings in Euclidean space, after that, by injecting them into the tangent space of the reference point, we derive the initial hyperbolic embeddings of both users and items:
(6) $z_{u}^{0}=(0,e_{u}),\quad\qquad z_{i}^{0}$ $=(0,e_{i})$
$h_{u}^{0}=\exp_{\mathbf{o}}(z_{u}^{0}),\quad\quad h_{i}^{0}$ $=\exp_{\mathbf{o}}(z_{i}^{0})$
where $e_{u},e_{i}$ are Gaussian distributed initialized Euclidean embeddings, and $z_{u}^{0},z_{i}^{0}$ are corresponding embeddings in the tangent space of the origin point. $h_{u}^{0},h_{i}^{0}$ therefore stands for the resulting initialized hyperbolic embeddings.
Hyperbolic Message Passing. Message passing between each node is realized through hyperbolic neighborhood aggregation, which is used to extract explicit user-item interaction. Given user $u$, item $i$, as well as their neighborhood $\mathcal{N}_{u},\mathcal{N}_{i}$, their embeddings at layer $l$ in the tangent space are updated as follows:
(7) $z_{u}^{l}=z_{u}^{l-1}+\sum_{i\in\mathcal{N}_{u}}\frac{1}{|\mathcal{N}_{u}|}z_{i}^{l-1},\quad\quad z_{i}^{l}=z_{i}^{l-1}+\sum_{u\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}z_{u}^{l-1}$
where $|\mathcal{N}_{u}|$ and $|\mathcal{N}_{i}|$ represent the number of neighbors of $u$ and $i$, respectively. The final tangent space embedding is derived by summing up embeddings at each layer:
(8) $z_{u}=\sum_{l}z_{u}^{l},\quad\quad z_{i}=\sum_{l}z_{i}^{l}$
Since $z_{u}$ and $z_{i}$ are on the tangent space, we still need to project it into hyperbolic space to derive the hyperbolic embeddings:
(9) $h_{u}=\exp_{\mathbf{o}}(z_{u}),\quad\quad h_{i}=\exp_{\mathbf{o}}(z_{i})$
Hyperbolic Prediction Function. Through hyperbolic messaging passing, we’ve captured high-order relational information into hyperbolic embeddings. Hyperbolic distance is now used to define a prediction function, which is used to predict the possibility that a user would interact with an item:
(10) $p(u,i)=\frac{1}{d_{\mathcal{H}}(h_{u},h_{i})}$
Hyperbolic Margin Ranking Loss. Bayesian Personalized Ranking (BPR) loss has been widely applied in recommender systems (Rendle et al., 2012), aiming to pull together user-item pairs with recorded interactions while pushing apart those without. In our model, we adopt the hyperbolic margin ranking loss (Sun et al., 2021), designed to prevent the undesirable collapse seen in Euclidean space by adding an additional margin to the original BPR loss:
(11) $\ell_{margin}(u,i,j)=\max(d_{\mathcal{H}}^{2}(h_{u},h_{i})-d_{\mathcal{H}}^{2}(h_{u},h_{j})+m,0),$
where $d_{\mathcal{H}}$ represents hyperbolic distance, $m$ is the margin, $(u,i)$ is an observed interaction, and $(u,j)$ is a sampled negative pair.
Hyperbolic Informative Mechanism. We further integrate a geometric-aware mechanism (Yang et al., 2022a) to enhance the pull-push process, consisting of two parts: Hyperbolic-aware Margin Learning (HAML) and Hyperbolic Informative Negative Sampling (HINS). Given that distances in hyperbolic space grow exponentially as embeddings deviate from the origin, we aim to assign a larger margin when user-item pairs are near the origin. HAML addresses this issue by dynamically adapting the margin as follows:
(12) $m^{\mathcal{H}}_{ui}$ $=\mathrm{sigmoid}(\delta),$
$\delta$ $=\frac{d_{\mathcal{H}}^{2}(e_{u},\mathbf{o})+d_{\mathcal{H}}^{2}(e_{i},\mathbf{o})-d_{\mathcal{H}}^{2}(e_{u},e_{i})}{e_{u,0}\cdot e_{i,0}}$
where $e_{u,0}$ and $e_{i,0}$ represent the zeroth coordinate elements of $e_{u}$ and $e_{i}$, respectively, corresponding to the embeddings norms in the hyperbolic space. HINS is an algorithm that samples negative nodes close to the positive item, yielding informative negative samples for both head and tail items. This is achieved by iteratively calculating the distances between the positive item and all candidate negative samples until the nearest one is identified.
4.2. Semantic Representations Generation
In this subsection, we introduce a paradigm for distilling useful information from raw text descriptions of items and users into dense semantic embeddings, which will be used later in our alignment framework that derives the final embeddings, integrating both collaborative information and semantic insights. Inspired by prior research (Ma et al., 2024; Ren et al., 2024) that leverages the power of large language models (LLMs) to extract useful user/item descriptions from raw, noisy textual data, we adopt a similar framework to generate user/item profiles with rich textual descriptions:
(13) $\mathcal{P}_{i}=\textit{LLMs}(\mathcal{S}_{i},\mathcal{T}_{i})$
where $\mathcal{S}_{i}$ represents system prompts to the LLMs, $\mathcal{T}_{i}$ denotes raw textual information, and $\mathcal{P}_{i}$ is the resulting user/item profile, consisting of sentence-level descriptions.
While we’ve derived textual information for users/items, it cannot be directly fused with the embeddings generated from hyperbolic collaborative filtering. To integrate both types of knowledge, we encode the textual information into embeddings using pretrained text encoder models, such as BERT (Devlin, 2018):
(14) $e_{i}=\textit{Encoder}(\mathcal{P}_{i})$
4.3. Hyperbolic Alignment
Given two sets of embeddings: hyperbolic embeddings with collaborative information and semantic embeddings with textual information, we aim to maximize the mutual information between them, creating unified embeddings that integrate both types of knowledge. To achieve this, we introduce a hyperbolic alignment loss, which mitigates the effects of noisy signals in representation learning and improves recommendations for tail items by aligning information within the same hyperbolic space. A theoretical proof of the advantages of alignment in hyperbolic space over Euclidean space is provided in the next subsection.
Before calculating the alignment loss, two steps are necessary: 1) the semantic embeddings must be adjusted to match the dimensionality of the collaborative embeddings, and 2) they should be transformed into a common representation space. To accomplish this, we first apply a multi-layer perceptron (MLP) to the semantic embeddings to ensure dimensional alignment. We then project the semantic embeddings into hyperbolic space so they can align with the hyperbolic collaborative embeddings:
(15) $s_{i}$ $=\exp_{x}(e_{i})$
$s^{\prime}_{i}$ $=\textit{MLP}(s_{i})$
where $\exp_{x}$ is defined as in 3, $e_{i}$ represents the semantic embeddings from the encoder, $s_{i}$ is the dimension-aligned semantic embedding, and $s^{\prime}_{i}$ denotes the final projected semantic embedding in hyperbolic space. The hyperbolic alignment loss is then calculated as:
(16) $\ell_{align}(i)=d_{\mathcal{H}}^{2}(h_{i},\textbf{s}^{\prime}_{i})$
where $d_{\mathcal{H}}$ represents the hyperbolic distance, and $h_{i}$ is the hyperbolic embedding with collaborative information.
4.4. Gradient Analysis of Semantic Alignment.
To further explore the impact of aligning semantic embeddings within the hyperbolic model, we analyze the gradient behavior with respect to nodes of varying norms in Proposition 4.1. This provides insight into the hierarchical preservation properties of the hyperbolic space, which are absent in Euclidean space.
Proposition 4.1.
The gradient magnitude of semantic alignment in hyperbolic space $\|\nabla_{\mathbf{x}}d_{\mathcal{H}}\|\approx\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}$ enables adaptive updates for hierarchical preservation, while Euclidean space maintains constant magnitude $\|\nabla_{\mathbf{x}}d_{E}\|=1$, where the $d_{\mathcal{H}}$ and $d_{\mathcal{E}}$ represent the hyperbolic distance and Euclidean distance, respectively.
Hyperbolic Space. Suppose $\mathbf{x}$ is a hyperbolic embedding derived from  4.1, and $\mathbf{y}$ is the semantic embedding from  4.2. The hyperbolic distance is given by:
(17) $d_{\mathcal{H}}(\mathbf{x},\mathbf{y})=\operatorname{arccosh}(z)$
where $z=-\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}$. For large $\|\mathbf{x}\|$ and $\|\mathbf{y}\|$, given $x_{0}=\sqrt{1+\|\mathbf{x}\|^{2}}$ and $y_{0}=\sqrt{1+\|\mathbf{y}\|^{2}}$, we adopt the following approximations for further calculation:
(18) $x_{0}\approx\|\mathbf{x}\|,\quad y_{0}\approx\|\mathbf{y}\|,\quad\sqrt{z^{2}-1}\approx z$
(19) $z=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}\approx\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)$
Let $x_{j}=\|\mathbf{x}\|\hat{x}_{j}$ and $y_{j}=\|\mathbf{y}\|\hat{y}_{j}$ (where $\hat{x}_{j}$ represents the direction of $x_{j}$). The partial derivative of $z$ with respect to $x_{j}$ is:
(20) $\frac{\partial z}{\partial x_{j}}=\frac{x_{j}}{x_{0}}y_{0}-y_{j}\approx\frac{x_{j}}{\|\mathbf{x}\|}\|\mathbf{y}\|-y_{j}=\|\mathbf{y}\|(\hat{x}_{j}-\hat{y}_{j})$
Thus, the gradient norm of $d_{\mathcal{H}}$ for $\mathbf{x}$ can be approximated by:
(21) $\|\nabla_{\mathbf{x}}d_{\mathcal{H}}\|=\frac{\|\nabla_{\mathbf{x}}z\|}{\sqrt{z^{2}-1}}\approx\frac{\|\nabla_{\mathbf{x}}z\|}{z}=\frac{\|\mathbf{y}\|\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)}=\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}$
This result indicates that nodes with large norms undergo smaller gradient updates, preserving local structures, while nodes with smaller norms adjust more, refining global hierarchical relationships and capturing hierarchical structure effectively.
Euclidean Space. For comparison, we examine the gradient behavior in Euclidean space. The Euclidean distance is defined as:
(22) $f(\mathbf{x})=d_{E}(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|$
The gradient of $f(\mathbf{x})$ with respect to $\mathbf{x}$ is:
(23) $\nabla_{\mathbf{x}}f(\mathbf{x})=\frac{\mathbf{x}-\mathbf{y}}{\|\mathbf{x}-\mathbf{y}\|}=\frac{\mathbf{x}-\mathbf{y}}{d_{E}(\mathbf{x},\mathbf{y})}$
which has a magnitude of:
(24) $\|\nabla_{\mathbf{x}}f(\mathbf{x})\|=1$
In summary, the analysis shows that hyperbolic space provides adaptive gradient magnitudes that depend on the norms of the nodes, effectively capturing and preserving hierarchical structures. In contrast, Euclidean space, with its constant gradient magnitude, cannot adapt to hierarchical nuances, making it less suitable for modeling hierarchical data in recommender systems.
4.5. Hyperbolic Hierarchical Structure
Figure 3. Hierarchical Structure Visualization.
In this subsection, we introduce a mechanism that allows users to control the level of exploration in their recommendations by constructing a hierarchy tree based on the learned hyperbolic embeddings, an example is shown in Figure 3. These embeddings integrate both collaborative information and language semantics, enabling more nuanced recommendations.
Hierarchy Tree. The primary motivation for constructing a hierarchy tree is to uncover hidden structures in the data that are not explicitly reflected in the dataset. For instance, consider the Amazon-CD dataset: a CD might be categorized under genres like “folk music”, but it cannot be labeled as “folk music preferred by teenagers” or further refined to “folk music preferred by teenagers who live in large cities”. However, in reality, such nuanced preferences exist within the collaborative information. When recommending based solely on past interactions, we may overlook these hidden layers of structure that are implicit in the dataset (e.g., “music” → “folk music” → “folk music preferred by teenagers” → “folk music preferred by teenagers who live in large cities”).
As previously discussed, in hyperbolic space, the norm of the embeddings reflects the level of exploration, where depth search facilitates exploitation and breadth search enables exploration. Building on this concept, we use the learned embeddings to form the bottom layer of the hierarchy tree. We then apply K-Means clustering in hyperbolic space to derive higher-level cluster nodes, which act as pseudo clustering nodes within the hierarchy, representing group preferences. For example, if an item is classified as “folk music preferred by teenagers who live in large cities”, its parent cluster may represent “folk music preferred by teenagers”, while its sibling clusters could include “folk music preferred by teenagers who live in small towns”. This hierarchical clustering process continues iteratively until reaching the root, thereby creating a multi-level tree structure that captures different levels of exploration.
Exploration-Exploitation Balance. To allow users control over the balance of exploration and exploitation, we introduce two adjustable parameters: temperature $\tau$ and hierarchy level $l$. In the standard recommendation approach, items are recommended based on similarity to the user’s embedding, which may limit exploration. For users with focused preferences, this method may consistently recommend items that align too closely with their past choices, offering little opportunity to explore different categories.
To address this issue, we modify the recommendation process by replacing a portion of the recommended items with alternatives from other branches of the hierarchy tree. This modification gives users control over the balance between exploration and exploitation. Temperature $\tau$ allows users to specify the proportion of original recommendations to replace, while level $l$ determines the extent of exploration within the hierarchy. For example, if a user selects $\tau=0.5$ and $l=2$, half of the originally recommended items will be retained, while the other half will be replaced by items sampled from nodes two levels up in the hierarchy tree, with selections made from the connected bottom-layer nodes.
5. EXPERIMENTS
Table 1. Comprehensive model comparison based on utility and diversity metrics, where higher values signify superior performance. The best-performing results for each metric are highlighted in bold, with the second-best results underlined.
Data Models utility $\uparrow$ Diveristy $\uparrow$
Recall@10 NDCG@10 Recall@20 NDCG@20 Div@10 H@10 EPC@10 Div@20 H@20 EPC@20
Amazon- books LightGCN 0.0950 0.0726 0.1428 0.0884 0.4469 9.8554 0.7429 0.4504 10.485 0.7798
TAG-CF 0.0986 0.0755 0.1468 0.0912 0.4090 10.915 0.7899 0.4517 11.465 0.8227
LightGCL 0.1075 0.0846 0.1515 0.0991 0.4651 11.511 0.8523 0.4974 11.941 0.8728
SimGCL 0.0997 0.0762 0.1478 0.0918 0.4286 10.341 0.7757 0.4802 10.915 0.8076
HCCF 0.1045 0.0813 0.1488 0.0962 0.5008 11.814 0.8579 0.5433 12.137 0.8746
HGCF 0.1013 0.0759 0.1511 0.0921 0.4742 11.577 0.8623 0.5041 11.732 0.8585
HICF 0.1113 0.0862 0.1599 0.1022 0.4702 11.922 0.8561 0.5261 12.129 0.8660
Ours 0.1167 0.0902 0.1682 0.1069 0.5255 12.096 0.8700 0.5754 12.254 0.8758
Yelp LightGCN 0.0678 0.0565 0.1135 0.0717 0.2603 8.3263 0.6842 0.2277 9.1811 0.7343
TAG-CF 0.0692 0.0571 0.1142 0.0722 0.2353 8.4516 0.6939 0.2276 9.3108 0.7428
LightGCL 0.0709 0.0594 0.1141 0.0739 0.2856 11.243 0.8281 0.3064 11.823 0.8550
SimGCL 0.0707 0.0586 0.1184 0.0748 0.3397 9.8191 0.7642 0.3460 10.508 0.7973
HCCF 0.0680 0.0562 0.1097 0.0701 0.3168 11.942 0.8572 0.3535 12.311 0.8752
HGCF 0.0654 0.0510 0.1099 0.0665 0.3133 11.977 0.8718 0.3289 12.103 0.8747
HICF 0.0709 0.0579 0.1169 0.0737 0.3478 11.877 0.8557 0.3721 12.147 0.8666
Ours 0.0739 0.0597 0.1224 0.0762 0.3874 12.155 0.8716 0.3961 12.329 0.8773
Google- reviews LightGCN 0.0932 0.0745 0.1430 0.0929 0.2422 9.9149 0.7241 0.2725 10.761 0.7749
TAG-CF 0.0942 0.0754 0.1447 0.0940 0.2376 10.184 0.7485 0.2497 11.031 0.7961
LightGCL 0.0945 0.0768 0.1461 0.0959 0.2639 11.643 0.8333 0.2705 12.317 0.8655
SimGCL 0.0979 0.0784 0.1482 0.0971 0.2966 10.361 0.7486 0.2950 11.119 0.7923
HCCF 0.0923 0.0744 0.1446 0.0937 0.2195 10.741 0.8056 0.2402 11.561 0.8433
HGCF 0.0910 0.0697 0.1485 0.0909 0.2859 12.248 0.8538 0.3321 12.474 0.8619
HICF 0.0951 0.0736 0.1515 0.0944 0.3262 12.376 0.8523 0.3510 12.660 0.8640
Ours 0.1002 0.0772 0.1598 0.0992 0.3185 12.466 0.8598 0.3543 12.769 0.8728
Table 2. Comparison with feature-enhanced Euclidean baselines, with the best-performing results highlighted in bold.
Data Amazon-books Yelp Google-reviews
Metrics R@20 N@20 Div@20 H@20 EPC@20 R@20 N@20 Div@20 H@20 EPC@20 R@20 N@20 Div@20 H@20 EPC@20
LightGCN+ 0.1466 0.0901 0.4349 10.3915 0.7802 0.1156 0.0733 0.2697 9.6903 0.7526 0.1438 0.0940 0.2707 10.8664 0.7801
LightGCL+ 0.1545 0.0992 0.5064 11.9761 0.8728 0.1131 0.0739 0.2759 11.8617 0.8581 0.1443 0.0946 0.2744 12.4700 0.8776
SimGCL+ 0.1494 0.0926 0.5163 11.4751 0.8342 0.1195 0.0752 0.3152 9.9952 0.7811 0.1504 0.0982 0.3030 11.2111 0.7953
HCCF+ 0.1525 0.0940 0.4316 11.7984 0.8616 0.1170 0.0729 0.2884 11.6906 0.8577 0.1438 0.0941 0.2524 11.5433 0.8406
Ours 0.1682 0.1069 0.5754 12.2543 0.8758 0.1224 0.0762 0.3961 12.3290 0.8773 0.1598 0.0992 0.3543 12.7690 0.8728
5.1. Experimental Settings
Datasets. To evaluate our proposed model, we utilize three widely-used public datasets: Amazon-books, which contains user purchase behaviors within the book category on Amazon; Yelp, which records customer ratings and reviews for restaurants on Yelp; and Google-reviews, which includes user reviews and business metadata from Google Maps. Detailed statistics for these datasets , including the distribution of head and tail items, are provided in Table 3. The high proportion of tail users (T80) across these datasets highlights the prevalence of long-tail distributions, supporting the assumption of a power-law distribution in real-world data.
Evaluation Metrics. We assess our recommended items using both utility and diversity metrics, demonstrating that our model achieves precise recommendations for users’ preference while enhancing diversity in the recommendation list.
For utility, we adopt two widely-used ranking metrics: Recall and NDCG, which measure model effectiveness. To evaluate diversity, we employ three metrics: Distance Diversity (Ziegler et al., 2005; Zhang and Hurley, 2008), Shannon Entropy (Mehta et al., 2012), and Expected Popularity Complement (Vargas and Castells, 2011).
Distance Diversity (Div) quantifies the variation among recommended items, encouraging diversity within the recommendation:
(25) $Div(u)=\frac{\sum_{i\in R_{u}}\sum_{j\in R_{u},j\neq i}d(i,j)}{\text{Number of}(i,j)\text{ pairs}}$
where $R_{u}$ is the recommendation list for user $u$, and $d(i,j)$ represents the distance between items $i$ and $j$. Number of $(i,j)$ pairs is calculated as $\binom{|R_{u}|}{2}=\frac{|R_{u}|(|R_{u}|-1)}{2}$. For fair comparison, we compute distances within the same representation space across models.
Shannon Entropy (H) quantifies the unpredictability within the recommendations, with higher entropy signifying greater diversity:
(26) $H=-\sum_{i\in\text{set}(\sum_{u}R_{u})}p(i)\log p(i)$
where $\sum_{u}R_{u}$ denotes all recommended items across users (including duplicates), and $p(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}R_{u})}\text{count}(i)$ is the proportion of item $i$ among all recommended items.
Expected Popularity Complement (EPC) assesses the extent of bias towards popular items by promoting niche recommendations:
(27) $EPC=1-\frac{1}{|\text{set}(\sum R_{u})|}\sum_{i\in\text{set}(\sum_{u}R_{u})}\frac{\text{pop}(i)}{\max_{i}(\text{pop}(i))}$
where $N_{u}$ is the set of items user $u$ has interacted with, and $\text{pop}(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}N_{u})}\text{count}(i)$ denotes item $i$’s popularity.
Baselines. We compare our model with state-of-the-art baselines in Euclidean and Hyperbolic spaces. Euclidean models include:
  • LightGCN (He et al., 2020): A lightweight graph recommender model by removing redundant neural modules in graph convolution.
  • TAG-CF (Ju et al., 2024): Enhances collaborative filtering by performing a single message-passing step only at inference time.
  • LightGCL (Cai et al., 2023): Uses singular value decomposition (SVD) for robust graph augmentation without data augmentations.
  • SimGCL (Yu et al., 2022): Replaces graph augmentations in contrastive learning by introducing random noise during initialization.
  • HCCF (Xia et al., 2022): Integrates hypergraph and contrastive learning, capturing both local and global user dependencies.
Hyperbolic GNN models include:
  • HGCF (Sun et al., 2021): Combines hypergraph learning with collaborative filtering using hyperbolic margin ranking loss.
  • HICF (Yang et al., 2022a): Enhances head and tail item performance with hyperbolic geometry, improving utility for long-tail items.
Implementation Details. We set embedding dimensions to 50 and a training batch size of 1024. Embeddings in hyperbolic space are optimized using Riemannian SGD (RSGD) (Bécigneul and Ganea, 2018; Sun et al., 2021), while the MLP adapter is optimized with Adam (Kingma, 2014). Training incorporates early stopping technique based on Recall@10 with 10 patience steps. For the hierarchical structure, we use a $1:8$ ratio for parent and child nodes. We generate user and item profiles with GPT-3.5-turbo and encode them with text-embedding-ada-002 (Neelakantan et al., 2022) for Amazon-books and Yelp, and BERT (Devlin, 2018) for Google-reviews. The profile decoder backbone is Llama 3.1-8B. All experiments are conducted on NVIDIA A40/A100 GPUs with 40GB memory.
Table 3. Statistics of the experimental data.
Dataset #User #Item #Interactions Density
All H20(%) T80(%)
[…]
5.2. Performance Comparison
Overall Comparison. The performance comparison across all models is presented in Table 1. Our HARec consistently outperforms the baselines in both utility and diversity metrics, achieving up to $5.49\%$ improvement in utility metrics and $11.39\%$ increase in diversity metrics. Notably, most baselines struggle to achieve strong performance in both utility and diversity simultaneously. For instance, while HICF performs second-best in utility metrics on the Amazon-books dataset, outperforming other baselines by a significant margin (excluding our model), it lacks the same advantage in diversity metrics, where several baselines achieve better results. This highlights the superiority of our model, which achieves state-of-the-art performance in both utility and diversity: a feat not previously attained by any baseline.
Baselines with Feature Enhancement. To ensure a fair comparison, we further enhance Euclidean baselines by integrating feature information. Specifically, we align semantic embeddings with the original Euclidean embeddings following the structure in Section 4.2, but in Euclidean space. The results are presented in Table 2, with enhanced baselines denoted by “+”. The data shows that our HARec still surpasses all baselines in both utility and diversity. Interestingly, while the feature-enhanced versions of the Euclidean baselines generally improve in utility metrics, they do not exhibit a stable improvement in diversity metrics. This aligns with our theoretical analysis in Section 4.5, which demonstrates the advantages of our alignment model in the hyperbolic space.
Table 4. Layer-wise norms within the hierarchy structure.
Data level-0 level-1 level-2 level-3 level-4
Amazon-books 0.2388 0.4859 0.5143 0.5889 0.6773
Yelp 0.1758 0.3500 0.4719 0.5590 0.6307
Google-reviews 0.4374 0.8585 1.0682 1.1731 1.2479
Figure 4. Ablation study on model variants.
Table 5. Performance on H20 (head) and T80 (tail) items, with the top hyperbolic results highlighted in bold. $\Delta_{\mathcal{H}}(\%)$ represents the improvement over the second-best hyperbolic baseline, which is underlined.
Data Amazon-books Yelp Google-reviews
Metrics Recall@20 NDCG@20 Recall@20 NDCG@20 Recall@20 NDCG@20
H20 T80 H20 T80 H20 T80 H20 T80 H20 T80 H20 T80
LightGCN 0.1162 0.0266 0.0744 0.0140 0.1076 0.0059 0.0693 0.0024 0.1311 0.0119 0.0878 0.0051
HGCF 0.1083 0.0428 0.0673 0.0248 0.0789 0.0310 0.0489 0.0176 0.1123 0.0362 0.0714 0.0195
HICF 0.1086 0.0513 0.0713 0.0309 0.0847 0.0322 0.0556 0.0180 0.1147 0.0388 0.0746 0.0208
Ours 0.1125 0.0557 0.0730 0.0339 0.0863 0.0361 0.0558 0.0204 0.1155 0.0443 0.0755 0.0238
$\Delta_{\mathcal{H}}(\%)$ +3.59 +8.58 +2.38 +9.71 +1.89 +12.11 +0.36 +13.33 +0.70 +14.18 +1.21 +14.42
Figure 5. Performance across different levels of exploration.
5.3. Ablation Study
We perform an ablation study to examine the contributions of two key components in our model: hyperbolic margin ranking loss and semantic alignment loss. To maintain meaningful evaluation, we focus solely on utility metrics, as random recommendations may artificially inflate diversity without practical utility. The following notations are used for our ablation variants: “w/o M” denotes the model without hyperbolic margin ranking loss, “w/o S” represents the model without semantic alignment loss, and “w/o M & S” indicates that both margin ranking and semantic alignment are excluded, leaving only the BPR loss.
Figure 4 shows that our full model outperforms all ablation variants, highlighting its overall effectiveness. Models lacking both margin ranking and semantic alignment perform the worst, while removing either module results in decreased utility, emphasizing the importance of each component and their effective integration.
5.4. Hierarchical Structure Analysis
To validate the effectiveness of our hierarchical structure, which allows users to control their level of exploration within recommendations, we conducted a series of analyses. Our goal is to demonstrate how the hierarchical model provides a structured balance between exploration and exploitation, allowing users to adjust the generality of recommendations by themselves.
First, we analyze the norms of node embeddings across hierarchical layers. As shown in Table 4, nodes in higher layers have consistently smaller norms. This observation supports our design intention: higher-layer nodes represent broader, generalized group preferences, aligning with findings in Section 4.5. This structure enables more abstract grouping as the layers ascend, thereby capturing progressively broader user interests.
To further examine the hierarchy’s practical utility, we designed experiments that simulate varying user preferences in daily interactions. Using a temperature parameter $\tau=0.2$ to control recommendation randomness, we tested recommendation performance across different hierarchical levels, from layer 00 (original recommendation level) up to layer $3$. The results, visualized in Figure 5, reveal a clear pattern: as the level $l$ increases, utility metrics tend to decrease, while diversity metrics increase.
This progression highlights the intended trade-off our model offers: recommendations from lower layers (e.g., layer 00) focus on utility and specific user interests, ideal for exploitation. In contrast, higher layers (e.g., layer $3$) prioritize diversity by presenting broader, less tailored options, thus encouraging exploration. This systematic change across levels provides users with a flexible mechanism to adjust recommendation focus: from niche, highly relevant items to a more varied, exploratory set—based on their current preference for exploration or exploitation.
To underscore the effectiveness of semantic alignment, particularly for cold-start (tail) items, we conducted experiments measuring utility performance on both head and tail items. In our context, head items (H20) refer to the top $20\%$ of items with the highest number of interactions in the training dataset, representing the most popular items. The remaining items are categorized as tail items (T80), which have limited interaction history, representing less frequently engaged (cold-start) items.
For simplicity, we compare our model with all hyperbolic baselines and a representative Euclidean baseline (LightGCN). The results, displayed in Table 5, reveal significant improvements, especially for tail items. Compared with hyperbolic baselines without semantic alignment, our model shows a notably larger performance boost on tail items than on head items. This finding indicates that semantic alignment particularly enhances the recommendation of cold-start items, effectively mitigating the cold-start problem for items with limited or no prior interactions.
6. CONCLUSION
In this work, we introduce HARec, a novel hyperbolic graph-text alignment framework that effectively balances exploration and exploitation by integrating semantic and collaborative information, with theoretical analysis supporting the advantages of this approach within hyperbolic space. Additionally, we propose a unique mechanism that enables personalized control over the exploration-exploitation trade-off through a hierarchical tree structure, allowing users to adjust recommendation preferences dynamically. Extensive experiments demonstrate that HARec achieves state-of-the-art performance, with up to a $5.49\%$ improvement in utility metrics and an $11.39\%$ increase in diversity metrics, making it the first model to excel in both criteria simultaneously, a significant achievement that has yet to be attained by any previous models.
Appendix A Appendix
In the supplementary materials, we provide additional details on our datasets and present several case studies illustrating the performance of our decoder module.
A.1. Profile Decoder
To further evaluate the expressiveness of our learned hyperbolic embeddings, we introduce a profile decoder designed to reconstruct the original profile from these embeddings, thereby broadening their application beyond recommendation tasks. Our method begins by mapping the hyperbolic embeddings back into Euclidean space to facilitate alignment with traditional network layers. These Euclidean-mapped embeddings are then processed through a mixture of experts (MoE) module. The MoE serves to refine the embeddings, ensuring consistency in both dimensionality and semantic representation (Ma et al., 2024). Once processed, these embeddings are fed into the language model, specifically after the positional layer, to incorporate them within a sequential architecture. To enhance training efficiency and minimize computational cost, we freeze the language model’s parameters, making the MoE module the only trainable component in the architecture.
Details of the input prompt structure and processing flow are depicted in Figure 6. To facilitate seamless embedding integration, we introduce two additional tokens, ¡USER_EMBED¿ and ¡ITEM_EMBED¿, to the language model’s tokenizer. This addition enables these embeddings to be treated as single, coherent tokens, ensuring accurate tokenization and embedding interpretation. After passing through the positional layer, the placeholders are dynamically replaced by the adapted embeddings generated by the MoE, allowing the embeddings to retain context-specific information.
A.2. Decoder Quality
To assess the effectiveness of our decoder mechanism, we compare the embeddings from our model against those from Euclidean embeddings (LightGCN) and Hyperbolic embeddings (HICF) as inputs to the decoder. The results are shown in Table 6. We use BERTScore (Zhang et al., [n. d.]) to evaluate the semantic alignment of generated sentences with the original profile. Results indicate that our HARec outperforms both Euclidean and hyperbolic baselines, demonstrating that our model’s embeddings support recommendation and retain user/item profile information. This finding underscores that our model captures collaborative information alongside recoverable language semantics: a capability not achieved by previous models.
A.2.1. Decoder Case Study
To vividly illustrate the semantic capacity of our embeddings, we present two case studies demonstrating the model’s ability to recover user and item profiles.
————————————————————————————————————— Original user profile:
Based on the user’s interactions and reviews, this user is likely to enjoy unique dining experiences with a variety of options, trendy atmosphere with good drink and food selections, premium quality frozen yogurt with a wide variety of toppings, creative burgers and hot dogs, delicious Thai cuisine, and authentic and spicy Chinese cuisine served family-style.
Recovered user profile:
Based on the user’s interactions, it is likely that they would enjoy businesses that offer a variety of cuisines, including Mexican, Italian, and Asian fusion. They appreciate good food, drinks, and a relaxed atmosphere. They also value quality and freshness in their food.
————————————————————————————————————— In this example, the profile recovered by our decoder accurately captures the user’s preferences for Thai and Chinese cuisines, broadly summarized as an affinity for Asian fusion flavours. Additionally, our decoder identifies the user’s interest in diverse culinary experiences, indicating a propensity toward exploration.
————————————————————————————————————— Original item profile:
Ceol Irish Pub in Reno is a popular bar and pub known for its happy hour deals, great service, and friendly atmosphere. It is suitable for those who enjoy trivia nights, live music, and a wide selection of drinks, particularly whiskey. The pub has a welcoming environment, with bartenders who are attentive and quick to serve.
Recovered item profile:
Users who enjoy a variety of beer options, live music, and a relaxed atmosphere would enjoy Ceol Irish Pub.
————————————————————————————————————— The alignment between the original and recovered item profiles shows that our decoder effectively captures key thematic elements of Ceol Irish Pub. The original profile emphasizes aspects such as happy hour, a wide selection of drinks, live music, and a welcoming atmosphere with attentive service. The recovered profile distills these characteristics, focusing on the pub’s appeal for users who enjoy diverse beer options, live music, and a relaxed setting. This alignment suggests that our decoder reliably summarizes key attributes, albeit with a degree of abstraction.
Figure 6. A depiction of model prompt instruction.
Table 6. Experiments assessing decoder quality.
Data Amazon-books Yelp
BERTScore P R F1 P R F1
LightGCN 0.3266 0.2854 0.3064 0.3511 0.2238 0.2872
HICF 0.3498 0.3009 0.3257 0.3608 0.2398 0.3021
Ours 0.3665 0.3167 0.3418 0.3716 0.2453 0.3081