### v3

# Breaking Information Cocoons: A Hyperbolic Graph-LLM Framework for Exploration and Exploitation in Recommender Systems

Abstract
Modern recommender systems often create information cocoons, restricting users’ exposure to diverse content. A key challenge lies in balancing content exploration and exploitation while allowing users to adjust their recommendation preferences. Intuitively, this balance can be modeled as a tree-structured representation, where depth search facilitates exploitation and breadth search enables exploration. However, existing approaches face two fundamental limitations: Euclidean methods struggle to capture hierarchical structures, while hyperbolic methods, despite their superior hierarchical modeling, lack semantic understanding of user and item profiles and fail to provide a principled mechanism for balancing exploration and exploitation. To address these challenges, we propose HERec, a hyperbolic graph-LLM framework that effectively balances exploration and exploitation in recommender systems. Our framework introduces two key innovations: (1) a semantic-enhanced hierarchical mechanism that aligns rich textual descriptions processed by large language models (LLMs) with collaborative information directly in hyperbolic space, allowing for more nuanced updates that respect the underlying hierarchical structure in user-item profiles; (2) an automatic hierarchical representation by optimizing Dasgupta’s cost, which discovers hierarchical structures without requiring predefined hyperparameters, enabling user-adjustable exploration-exploitation trade-offs. Extensive experiments demonstrate that HERec consistently outperforms both Euclidean and hyperbolic baselines, achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics, effectively mitigating information cocoons.
1 INTRODUCTION
As online platforms grow in size dramatically, users are overwhelmed by an influx of information, creating information cocoons that restrict their exposure to diverse and novel content [16, 4].
Traditional recommendation approaches have primarily relied on collaborative filtering and interaction data [11, 14], where users with similar historical preferences are expected to share future interests [6]. However, these methods face two critical challenges in improving user experience. First, existing models struggle to capture the underlying hierarchical structure inherent in user-item networks [48, 27], a critical factor for enhancing performance and personalization. Second, these models fail to effectively balance the exploration-exploitation trade-off. For instance, a jazz enthusiast may want to explore rock music, but current models primarily focus on exploitation—recommending only familiar genres—without adequately supporting exploration.
Hyperbolic space has emerged as a promising approach for modeling hierarchical structures and power-law distributions [15, 37, 25, 36], where its exponential volume growth naturally aligns with hierarchical, scale-free structures [10, 5, 18, 46, 38]. In hyperbolic space, distances increase exponentially from the origin, allowing popular items or exploratory users to be positioned closer to the origin, while niche items or users with focused interests spread toward the boundaries. This spatial configuration enables hyperbolic graph neural network (GNN) models to perceive hierarchical relationships effectively, enhancing recommendation performance [36, 25, 37].
Despite these advantages, current hyperbolic recommender systems still face key limitations. First, they struggle with a lack of semantic understanding of user and item profiles. While large language models (LLMs) offer a potential solution, their text encoders operate in Euclidean space, making it challenging to integrate LLM-derived semantic features with hyperbolic collaborative signals. This mismatch leaves existing hyperbolic models vulnerable to noise in collaborative information, as they fail to capture semantic insights effectively. Second, hyperbolic recommender systems lack a principled approach to balancing exploration (introducing new content) and exploitation (focusing on existing preferences) [32]. This limitation reduces their ability to adapt to the exploration-exploitation trade-off, ultimately diminishing both the utility and diversity of recommendations.
Proposed Method. To address these challenges, we propose HERec, a hyperbolic graph-llm framework designed to balance exploration and exploitation in recommender systems. Our approach introduces two key innovations: (1) a hierarchical-aware graph-LLM alignment mechanism, which jointly aligns textual descriptions processed by LLMs with user-item collaborative information in hyperbolic space, supported by theoretical gradient analysis demonstrating its adaptiveness; (2) a novel hierarchical representation structure that enables user-adjustable exploration-exploitation trade-offs. Unlike conventional approaches that require manually defined hierarchies, we propose a hyperparameter-free clustering mechanism theoretically optimized by Dasgupta’s cost [7]. This approach automatically discovers hierarchical structures without requiring predefined hyperparameters, allowing for a more adaptive and principled organization of user preferences.
To validate the effectiveness of HERec, we conducted extensive experiments across multiple dimensions. Our model consistently outperformed baselines in both utility and diversity, marking a significant achievement as it is the first model to excel in both aspects simultaneously. For a fair comparison, we also enhanced Euclidean baselines with feature information, yet our model still demonstrated superior performance. Additionally, we conducted a detailed analysis on head and tail items, showing that HERec effectively boosts recommendations for tail items, thus enhancing diversity. An ablation study, along with a hierarchical representation structure analysis module, confirms HERec’s capability to leverage both semantic meaning and collaborative information, providing deeper insights into the advantages of our approach.
Contributions. Our primary contributions focus on three key aspects: capturing collaborative signals, integrating semantic information, and modeling the underlying hierarchical structure. The proposed graph-llm alignment framework unifies collaborative and semantic information in hyperbolic space, while a hierarchical representation structure organizes user preferences, allowing for dynamic adjustment between exploration and exploitation. Moreover, through extensive experiments evaluating both utility and diversity metrics, we demonstrate that HERec achieves state-of-the-art performance across utility and diversity metrics, making it the first model to excel in both criteria simultaneously. We have made our model implementation available at: https://github.com/Martin-qyma/HERec.
2 Related Work
In this section, we review graph-based collaborative filtering, large language models (LLMs) for recommendation, and hyperbolic representation learning methods.
Graph Collaborative Filtering. Graph Collaborative Filtering has emerged as a dominant approach in recommender systems, with graph neural networks (GNNs) playing a central role. Unlike traditional collaborative filtering techniques, such as matrix factorization [14], GNNs capitalize on the relational structure in user-item interactions, enabling a more nuanced understanding of complex dependencies. Prominent graph-based methods, such as NGCF [30], SGL [33], and LightGCN [11], have been widely adopted and have consistently achieved state-of-the-art performance. Nevertheless, these models are constrained to Euclidean space, which may fall short in capturing the underlying power-law distribution in user-item networks.
Large Language Models for Recommendation. Large language models (LLMs) are increasingly used in recommender systems for their strong capability in understanding textual information. Some approaches leverage LLMs directly for downstream recommendation tasks [1, 17, 19, 47, 45], while others utilize LLMs as an auxiliary tool for collaborative filtering [23, 34, 31]. However, these approaches require substantial inference time and are computationally inefficient. In our HERec, we use language models exclusively to improve data quality, eliminating the need for repeated inference during each recommendation by utilizing hyperbolic embeddings for the final output.
Hyperbolic Representation Learning. To address the limitations of Euclidean space in modeling hierarchical structures, researchers have investigated the use of hyperbolic space [5, 18, 37, 46], where distances grow exponentially as embeddings move away from the origin. This property makes hyperbolic geometry particularly well-suited for modeling user-item interactions by more accurately capturing latent hierarchies and dependencies. Consequently, hyperbolic representation learning has gained considerable attention in the recommender systems domain [41, 29, 25, 36], demonstrating improved performance over traditional Euclidean models.
However, none of these hyperbolic models fully leverage the rich semantic information embedded in language-based datasets. While some approaches incorporate knowledge graphs as supplementary information [26, 9], these models primarily focus on refining hierarchical structures rather than achieving a deeper semantic understanding of each item or user.
3 PRELIMINARIES
Embeddings learned from graph collaborative filtering are by default in the Euclidean space, which is unable to capture hierarchical structures. To address the problem, we introduce how to derive these embeddings in a hyperbolic space.
Riemannian Manifold. A Riemannian manifold $(\mathcal{M},g)$ is a smooth manifold with an inner product $g_{x}$ at each point $x\in\mathcal{M}$, defining geometric properties such as angles and curve lengths. Its curvature determines the geometry: positive (elliptic), zero (Euclidean), or negative (hyperbolic). We focus on hyperbolic geometry, modeled using the Lorentz (hyperboloid) representation.
Lorentz Model. An $n$-dimensional Lorentz manifold with negative curvature $-1/\kappa$ $(\kappa>0)$ is described as the Riemannian manifold $(\mathbb{H}^{n}_{\kappa},g_{\mathcal{L}})$. Here, $\mathbb{H}^{n}_{\kappa}=\{x\in\mathbb{R}^{n+1}:\langle x,x\rangle_{\mathcal{L}}=-\kappa,x_{0}>0\}$, $g_{\mathcal{L}}=\eta$ with $\eta=\mathbf{I}_{n}$ except $\eta_{0,0}=-1$, and $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ represents the Lorentzian inner product, which is defined as:
$\langle x,y\rangle_{\mathcal{L}}:=-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}$ (1)
A tangent space is an $n$-dimensional vector space that locally approximates hyperbolic space. For a point $x\in\mathbb{H}_{\kappa}^{n}$, the tangent space $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$ is defined as:
$\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}:=\{v\in\mathbb{R}^{n+1}:\langle v,x\rangle_{\mathcal{L}}=0\}$ (2)
This space is orthogonal to $x$ under the Lorentzian inner product, providing accurate local approximation. Transitions between hyperbolic space and tangent space are facilitated by exponential and logarithmic maps, essential for gradient-based optimization. The exponential map $\exp_{x}:\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}\to\mathbb{H}_{\kappa}^{n}$ projects a tangent vector onto hyperbolic space along a geodesic:
$\exp_{x}^{\kappa}(v)=\cosh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)x+\sqrt{\kappa}\sinh\left(\frac{\|v\|_{\mathcal{L}}}{\sqrt{\kappa}}\right)\frac{v}{\|v\|_{\mathcal{L}}}$ (3)
where $\|v\|_{\mathcal{L}}=\sqrt{\langle v,v\rangle_{\mathcal{L}}}$ is the Lorentzian norm. The logarithmic map $\log_{x}$ is the inverse of $\exp_{x}$ and retrieves the tangent vector for a point $y\in\mathbb{H}_{\kappa}^{n}$:
$\log_{x}^{\kappa}(y)=d_{\mathcal{L}}^{\kappa}(x,y)\frac{y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x}{\|y+\frac{1}{\kappa}\langle x,y\rangle_{\mathcal{L}}x\|}$ (4)
where $d_{\mathcal{H}}^{\kappa}(x,y)$ is the distance between $x$ and $y$ in $\mathbb{H}^{n}_{\kappa}$:
$d_{\mathcal{H}}^{\kappa}(x,y)=\sqrt{\kappa}\,\text{arcosh}\left(-\langle x,y\rangle_{\mathcal{L}}/\kappa\right)$ (5)
We use $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ as the reference point. For simplicity, we set $\kappa=1$, giving a curvature of $-1$, and omit $\kappa$ in subsequent discussions.
4 METHODOLOGY
In this section, we present a comprehensive overview of our proposed model, HERec. Our approach not only aims to deliver accurate recommendations but also addresses users’ growing demand for more diverse suggestions. Furthermore, by constructing a hierarchical tree from the learned embeddings, we introduce an additional mechanism that enables users to control the degree of exploration in the recommended items. The overall framework is in Figure 1.
Figure 1: The overall architecture of HERec. (i) Hyperbolic Graph Collaborative Filtering: Encodes collaborative information using hyperbolic GNNs; (ii) Hyperbolic Alignment: Aligns semantic and collaborative information within hyperbolic space; (iii) Hierarchical Representation Structure: Builds hierarchy structure from hyperbolic embeddings.
4.1 Hyperbolic Graph Collaborative Filtering
Similar to Euclidean space, hyperbolic graph collaborative filtering captures user-item dependencies via message passing mechanism. We apply the Lorentz representation for both users and items.
Hyperbolic embedding initialization. As discussed in section 3, we fix the origin $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ and use it as the reference point. Embeddings are initialized in Euclidean space using Gaussian sampling and projected onto the tangent space of $\mathbf{o}$ to derive their hyperbolic representations:
$z_{u}^{0}=(0,e_{u}),\quad\qquad z_{i}^{0}$ $=(0,e_{i})$ (6)
$h_{u}^{0}=\exp_{\mathbf{o}}(z_{u}^{0}),\quad\quad h_{i}^{0}$ $=\exp_{\mathbf{o}}(z_{i}^{0})$
where $e_{u},e_{i}$ are Gaussian-initialized Euclidean embeddings, and $z_{u}^{0},z_{i}^{0}$ are their tangent space projections. $h_{u}^{0},h_{i}^{0}$ denote the resulting hyperbolic embeddings.
Hyperbolic Message Passing. User-item interactions are captured via hyperbolic neighborhood aggregation. For user $u$, item $i$, and their neighborhoods $\mathcal{N}_{u},\mathcal{N}_{i}$, tangent space embeddings at layer $l$ are updated as:
$z_{u}^{l}=z_{u}^{l-1}+\sum_{i\in\mathcal{N}_{u}}\frac{1}{|\mathcal{N}_{u}|}z_{i}^{l-1},\quad z_{i}^{l}=z_{i}^{l-1}+\sum_{u\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}z_{u}^{l-1}$ (7)
where $|\mathcal{N}_{u}|$ and $|\mathcal{N}_{i}|$ are the neighborhood sizes. The final tangent embeddings are obtained by summing over all layers:
$z_{u}=\sum_{l}z_{u}^{l},\quad\quad z_{i}=\sum_{l}z_{i}^{l}$ (8)
Since $z_{u}$ and $z_{i}$ are on the tangent space, we still need to project them into hyperbolic space to derive the hyperbolic embeddings:
$h_{u}=\exp_{\mathbf{o}}(z_{u}),\quad\quad h_{i}=\exp_{\mathbf{o}}(z_{i})$ (9)
Hyperbolic Prediction Function. Through hyperbolic messaging passing, we’ve captured high-order relational information into hyperbolic embeddings. Hyperbolic distance is now used to define a prediction function, which predicts the possibility that a user would interact with an item:
$p(u,i)=\frac{1}{d_{\mathcal{H}}(h_{u},h_{i})}$ (10)
Hyperbolic Margin Ranking and Informative Mechanism. To address the limitations of Euclidean-based approaches, we adopt the hyperbolic margin ranking loss [25], an extension of Bayesian Personalized Ranking (BPR) [24]. This loss pulls user-item pairs with observed interactions closer while pushing negative pairs apart, preventing the collapse seen in Euclidean space by incorporating a margin:
$\ell_{m}(u,i,j)=\max(d_{\mathcal{H}}^{2}(h_{u},h_{i})-d_{\mathcal{H}}^{2}(h_{u},h_{j})+m,0)$ (11)
where $d_{\mathcal{H}}$ represents hyperbolic distance, $m$ is the margin, $(u,i)$ is an observed interaction, and $(u,j)$ is a sampled negative pair.
To further refine this pull-push process, we integrate a geometric-aware mechanism [36] combining Hyperbolic-aware Margin Learning (HAML) and Hyperbolic Informative Negative Sampling (HINS). HAML dynamically adjusts the margin based on hyperbolic geometry, assigning a larger margin to pairs near the origin:
$m^{\mathcal{H}}_{ui}$ $=\mathrm{sigmoid}(\delta)$ (12)
$\delta$ $=\frac{d_{\mathcal{H}}^{2}(e_{u},\mathbf{o})+d_{\mathcal{H}}^{2}(e_{i},\mathbf{o})-d_{\mathcal{H}}^{2}(e_{u},e_{i})}{e_{u,0}\cdot e_{i,0}}$
where $e_{u,0}$ and $e_{i,0}$ are the zeroth coordinates of $e_{u}$ and $e_{i}$, representing their norms. HINS enhances negative sampling by prioritizing candidates near the positive item, leveraging hyperbolic geometry to better model head and tail items, ultimately improving recommendation performance.
4.2 Semantic Representations Generation
We propose a paradigm to distill useful information from raw text descriptions of items and users into dense semantic embeddings for our alignment framework. Building on prior work [20, 23] that leverages large language models (LLMs) to extract meaningful user/item descriptions from noisy textual data, we generate rich user/item profiles as $\mathcal{P}_{i}=\textit{LLMs}(\mathcal{S}_{i},\mathcal{T}_{i})$, where $\mathcal{S}_{i}$ are system prompts, $\mathcal{T}_{i}$ is raw textual input, and $\mathcal{P}_{i}$ contains sentence-level descriptions.
To integrate these profiles with hyperbolic collaborative filtering embeddings, we encode the textual information into embeddings using pre-trained text encoders: $e_{i}=\textit{Encoder}(\mathcal{P}_{i})$. While LLM embeddings are an alternative, they often incur higher computational overhead and are less optimized for alignment with hyperbolic representations, making pre-trained encoders a more practical choice.
4.3 Hyperbolic Alignment
To maximize mutual information between hyperbolic embeddings (collaborative information) and semantic embeddings (textual information), we introduce a hyperbolic alignment framework to create unified representations that integrate both knowledge types. To achieve this, we develop a hyperbolic alignment loss, which reduces noise in representations and improves tail item recommendations by aligning information in the same hyperbolic space. A theoretical advantage of hyperbolic alignment over Euclidean alignment is discussed in the next subsection.
Before calculating the alignment loss, two steps are necessary: 1) the semantic embeddings must be adjusted to match the dimensionality of the collaborative embeddings, and 2) they should be transformed into a common representation space. To accomplish this, we first apply a multi-layer perceptron (MLP) to the semantic embeddings to ensure dimensional alignment. We then project the semantic embeddings into hyperbolic space so they can align with the hyperbolic collaborative embeddings. The hyperbolic alignment loss is calculated as:
$\ell_{align}(i)=d_{\mathcal{H}}^{2}(h_{i},s_{i})$ (13)
where $d_{\mathcal{H}}$ represents the hyperbolic distance, $h_{i}$ is the hyperbolic embedding with collaborative information and $s_{i}$ denotes the projected semantic embedding in hyperbolic space.
4.4 Gradient Analysis of Hyperbolic Alignment
To further explore the impact of aligning semantic embeddings within the hyperbolic model, we analyze the gradient behavior with respect to nodes of varying norms in Proposition 1, detailed proof with analysis can be found in Appendix A. This provides insight into the hierarchical preservation properties of the hyperbolic space, which are absent in Euclidean space.
Proposition 1.
Let $\mathbf{x},\mathbf{y}$ denote embeddings in hyperbolic space $\mathcal{H}^{n}$ with corresponding unit direction vectors in space-like dimension $\hat{\mathbf{x}},\hat{\mathbf{y}}$, and let $\theta$ be the angle between them. The gradient magnitude of semantic alignment satisfies:
$\|\nabla_{\mathbf{x}}d_{\mathcal{H}}(\mathbf{x},\mathbf{y})\|\approx\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}.$ (14)
This facilitates adaptive gradient updates that intrinsically preserve hierarchical structures. In contrast, Euclidean space exhibits constant gradient magnitude $\|\nabla_{\mathbf{x}}d_{\mathcal{E}}(\mathbf{x},\mathbf{y})\|=1$, where $d_{\mathcal{H}}$ and $d_{\mathcal{E}}$ denote hyperbolic and Euclidean distances respectively.
This proposition highlights that hyperbolic space offers adaptive gradient magnitudes based on node norms. Nodes with large norms (fine-grained preferences, small positional changes cause significant distance changes) receive smaller gradient updates, ensuring precise distance adjustments and preserving local structures. Meanwhile, nodes with smaller norms (abstract preferences, positional changes have less impact on distances) adjust more, refining global relationships and effectively capturing the hierarchical structure. This inherent ability of hyperbolic space to modulate learning intensity based on an embedding’s position within the hierarchy offers a distinct advantage over the fixed-update approach of Euclidean geometry, ultimately facilitating more faithful and structured representations.
4.5 Hierarchical Representation Structure
In this subsection, we design a mechanism that allows users to balance the exploration-exploitation trade-off by constructing a hierarchical representation structure based on the learned hyperbolic embeddings. A detailed complexity analysis is provided in Appendix C.2.
Hierarchy Tree. The primary motivation for constructing a hierarchy tree is to uncover hidden structures in the data that are not explicitly reflected in the dataset. For instance, consider the Amazon-CD dataset: a CD might be categorized under genres like “folk music”, but it cannot be labeled as “folk music preferred by teenagers”. However, in reality, such nuanced preferences exist within the collaborative information. When recommending based solely on past interactions, we may overlook these hidden layers of structure that are implicit in the dataset (e.g., “music” → “folk music” → “folk music preferred by teenagers”).
To address this, we introduce a hyperbolic hierarchical clustering method based on hyperbolic embeddings to build the hierarchy tree. Specifically, we iteratively group embeddings into clusters and use the centroid of each cluster as the data point to be clustered in the upper layer, referred to as a pseudo-cluster node. However, the number of cluster nodes in each layer is not predefined and must be determined. Inspired by Dasgupta’s cost [7], which encourages pushing edge cuts as far down the tree as possible, the optimal tree is required to be binary. Therefore, we fix the proportion of cluster nodes between layers as $k=2$. Details of the algorithm are presented in Appendix C.1.
Exploration-Exploitation Balance. To allow users to balance the exploration-exploitation trade-off, we introduce two adjustable parameters: temperature $\tau$ and hierarchy layer $l$. In the standard recommendation approach, items are recommended based on similarity to the user’s embedding, which may limit exploration. For users with focused preferences, this method may consistently recommend items that align too closely with their past choices, offering little opportunity to explore different categories.
To address this issue, we modify the recommendation process by replacing a portion of the recommended items with alternatives from other branches in the hierarchy tree. This modification gives users control over the balance between exploration and exploitation. Temperature $\tau$ allows users to specify the proportion of original recommendations to replace, while layer $l$ determines the starting point for exploration. For example, suppose a user selects $\tau=0.5$ and $l=5$. In that case, half of the original recommendations are retained, and the other half is replaced with items sampled from its ancestor cluster at layer $5$ of the hierarchy tree.
5 EXPERIMENTS
Table 1: Comprehensive model comparison based on utility and diversity metrics, where higher values signify superior performance. The best-performing results for each metric are highlighted in bold, with the second-best results underlined.
Data Models Utility $\uparrow$ Diveristy $\uparrow$
Recall@10 NDCG@10 Recall@20 NDCG@20 Div@10 H@10 EPC@10 Div@20 H@20 EPC@20
Amazon- books LightGCN 0.0950 0.0726 0.1428 0.0884 0.4469 9.8554 0.7429 0.4504 10.485 0.7798
TAG-CF 0.0986 0.0755 0.1468 0.0912 0.4090 10.915 0.7899 0.4517 11.465 0.8227
LightGCL 0.1075 0.0846 0.1515 0.0991 0.4651 11.511 0.8523 0.4974 11.941 0.8728
SimGCL 0.0997 0.0762 0.1478 0.0918 0.4286 10.341 0.7757 0.4802 10.915 0.8076
HCCF 0.1045 0.0813 0.1488 0.0962 0.5008 11.814 0.8579 0.5433 12.137 0.8746
HGCF 0.1013 0.0759 0.1511 0.0921 0.4742 11.577 0.8623 0.5041 11.732 0.8585
HICF 0.1113 0.0862 0.1599 0.1022 0.4702 11.922 0.8561 0.5261 12.129 0.8660
Ours 0.1167 0.0902 0.1682 0.1069 0.5255 12.096 0.8700 0.5754 12.254 0.8758
Yelp LightGCN 0.0678 0.0565 0.1135 0.0717 0.2603 8.3263 0.6842 0.2277 9.1811 0.7343
TAG-CF 0.0692 0.0571 0.1142 0.0722 0.2353 8.4516 0.6939 0.2276 9.3108 0.7428
LightGCL 0.0709 0.0594 0.1141 0.0739 0.2856 11.243 0.8281 0.3064 11.823 0.8550
SimGCL 0.0707 0.0586 0.1184 0.0748 0.3397 9.8191 0.7642 0.3460 10.508 0.7973
HCCF 0.0680 0.0562 0.1097 0.0701 0.3168 11.942 0.8572 0.3535 12.311 0.8752
HGCF 0.0654 0.0510 0.1099 0.0665 0.3133 11.977 0.8718 0.3289 12.103 0.8747
HICF 0.0709 0.0579 0.1169 0.0737 0.3478 11.877 0.8557 0.3721 12.147 0.8666
Ours 0.0739 0.0597 0.1224 0.0762 0.3874 12.155 0.8716 0.3961 12.329 0.8773
Google- reviews LightGCN 0.0932 0.0745 0.1430 0.0929 0.2422 9.9149 0.7241 0.2725 10.761 0.7749
TAG-CF 0.0942 0.0754 0.1447 0.0940 0.2376 10.184 0.7485 0.2497 11.031 0.7961
LightGCL 0.0945 0.0768 0.1461 0.0959 0.2639 11.643 0.8333 0.2705 12.317 0.8655
SimGCL 0.0979 0.0784 0.1482 0.0971 0.2966 10.361 0.7486 0.2950 11.119 0.7923
HCCF 0.0923 0.0744 0.1446 0.0937 0.2195 10.741 0.8056 0.2402 11.561 0.8433
HGCF 0.0910 0.0697 0.1485 0.0909 0.2859 12.248 0.8538 0.3321 12.474 0.8619
HICF 0.0951 0.0736 0.1515 0.0944 0.3262 12.376 0.8523 0.3510 12.660 0.8640
Ours 0.1002 0.0772 0.1598 0.0992 0.3185 12.466 0.8598 0.3543 12.769 0.8728
5.1 Experimental Settings
Datasets. To evaluate our proposed model, we utilize three widely-used public datasets: Amazon-books, which contains user purchase behaviors within the book category on Amazon; Yelp, which records customer ratings and reviews for restaurants on Yelp; and Google-reviews, which includes user reviews and business metadata from Google Maps. Detailed statistics, including the distribution of head and tail items, are provided in Appendix B. The high proportion of tail users (T80) highlights the prevalence of long-tail distributions, supporting the assumption of a power-law distribution in real-world data. More implementation details can be found in Appendix B.1.
Evaluation Metrics. We assess our recommended items using both utility and diversity metrics, demonstrating that our model achieves precise recommendations for users’ preference while enhancing diversity in the recommendation. For utility, we adopt two widely used ranking metrics: Recall and NDCG, which measure the effectiveness of the model. To evaluate diversity, we employ three metrics: Distance Diversity (Div) [49, 43], Shannon Entropy (H) [21], and Expected Popularity Complement (EPC) [28]. For more details, please refer to Appendix B.2.
Baselines. We compare our model with state-of-the-art Euclidean and Hyperbolic baselines:
  • LightGCN [11]: A lightweight graph convolution neural network model for recommendation by removing redundant neural modules in graph convolution.
  • TAG-CF [12]: A test-time augmentation framework that enhances recommendation by performing a single message-passing step only at inference time.
  • LightGCL [3]: A contrastive learning model that uses singular value decomposition (SVD) for robust graph augmentation without data augmentation.
  • SimGCL [42]: A contrastive learning model that replaces graph augmentations in contrastive learning by introducing random noise during initialization.
  • HCCF [35]: A collaborative filtering framework that integrates hypergraph and contrastive learning, capturing both local and global user dependencies.
  • HGCF [25]: A hyperbolic GNN model that combines hypergraph learning with collaborative filtering using hyperbolic margin ranking loss.
  • HICF [36]: A hyperbolic recommendation model that enhances head and tail item performance with hyperbolic geometry, improving utility for long-tail items.
5.2 Performance Comparison
Overall Comparison. The performance comparison across all models is presented in Table 1, with statistical significance testing in Appendix D.2. Our HERec consistently outperforms the baselines in both utility and diversity metrics, achieving up to $5.49\%$ improvement in utility metrics and $11.39\%$ increase in diversity metrics. Notably, most baselines struggle to achieve strong performance in both utility and diversity simultaneously. For instance, while HICF performs second-best in utility metrics on the Amazon-books dataset, it lacks the same advantage in diversity metrics, where several baselines achieve better results. This highlights the superiority of our model, which achieves state-of-the-art performance in both utility and diversity: a feat not previously attained by any baseline.
Baselines with Feature Enhancement. To ensure a fair comparison, we further enhance Euclidean baselines by integrating feature information. Specifically, we align semantic embeddings with the original Euclidean embeddings following the structure in Section 4.2, but in Euclidean space. The results are presented in Appendix D.1, with enhanced baselines denoted by “+”. The data shows that our HERec still surpasses all baselines in both utility and diversity. Interestingly, while the feature-enhanced versions of the Euclidean baselines generally improve in utility metrics, they do not exhibit a stable improvement in diversity metrics. This aligns with our theoretical analysis in Section 4.4, which demonstrates the advantages of our alignment in the hyperbolic space.
Profile Decoder. To validate the effectiveness of our alignment framework in capturing semantics, we designed a profile decoder module to reconstruct user and item profiles from the learned hyperbolic embeddings. The results demonstrate that profiles reconstructed from our model’s embeddings exhibit higher alignment with the originals, confirming the model’s ability to effectively capture underlying semantic meaning. Detailed methodology and experimental results are provided in Appendix E.
Performance on Head and Tail Items. To underscore the effectiveness of semantic alignment, particularly for cold-start (tail) items, we conducted experiments measuring utility performance on both head and tail items. In our context, head items (H20) refer to the top $20\%$ of items with the highest number of interactions in the training dataset, representing the most popular items. The remaining items are categorized as tail items (T80), which have limited interaction history, representing less frequently engaged items.
For simplicity, we compare our model with all hyperbolic baselines and a representative Euclidean baseline (LightGCN). The results, displayed in Table 2, reveal significant improvements, especially for tail items. Compared with hyperbolic baselines without semantic alignment, our model shows a notably larger performance boost on tail items than on head items. This finding indicates that semantic alignment particularly enhances the recommendation of cold-start items, effectively mitigating the cold-start problem for items with limited or no prior interactions.
Table 2: Performance on H20 (head) and T80 (tail) items, with the top hyperbolic results highlighted in bold. $\Delta_{\mathcal{H}}(\%)$ represents the improvement over the second-best hyperbolic baseline (underlined).
Datasets Amazon-books Yelp Google-reviews
Metrics Recall@20 NDCG@20 Recall@20 NDCG@20 Recall@20 NDCG@20
H20 T80 H20 T80 H20 T80 H20 T80 H20 T80 H20 T80
LightGCN 0.1162 0.0266 0.0744 0.0140 0.1076 0.0059 0.0693 0.0024 0.1311 0.0119 0.0878 0.0051
HGCF 0.1083 0.0428 0.0673 0.0248 0.0789 0.0310 0.0489 0.0176 0.1123 0.0362 0.0714 0.0195
HICF 0.1086 0.0513 0.0713 0.0309 0.0847 0.0322 0.0556 0.0180 0.1147 0.0388 0.0746 0.0208
Ours 0.1125 0.0557 0.0730 0.0339 0.0863 0.0361 0.0558 0.0204 0.1155 0.0443 0.0755 0.0238
$\Delta_{\mathcal{H}}(\%)$ +3.59 +8.58 +2.38 +9.71 +1.89 +12.11 +0.36 +13.33 +0.70 +14.18 +1.21 +14.42
Figure 2: Ablation study on model variants.
Figure 3: Layer-wise norms.
Figure 4: Performance analysis across hierarchical structure layers.
5.3 Ablation Study
We perform an ablation study to examine the contributions of two key components in our model: hyperbolic margin ranking loss and semantic alignment loss. To maintain meaningful evaluation, we focus solely on utility metrics, as random recommendations may artificially inflate diversity without practical utility. The following notations are used for our ablation variants: “w/o M” denotes the model without hyperbolic margin ranking loss, “w/o S” represents the model without semantic alignment loss, and “w/o M & S” indicates that both margin ranking and semantic alignment are excluded, leaving only the BPR loss. Figure 3 shows that our full model outperforms all ablation variants, highlighting its overall effectiveness. Models lacking both margin ranking and semantic alignment perform the worst, while removing either module results in decreased utility, emphasizing the effectiveness of integration.
Recent work [40, 39] has proposed performing message passing directly in hyperbolic space as an alternative to using the tangent space. We implemented this approach in place of our original hyperbolic graph collaborative filtering method. As shown in Appendix D.3, this alternative achieves comparable performance, making it a viable alternative.
5.4 Hierarchical Representation Structure Analysis
We evaluated the effectiveness of our hierarchical representation structure through experiments simulating varying user preferences. Using a temperature parameter $\tau=0.5$ to control recommendation randomness, we analyzed performance across various hierarchical layers. As shown in Figure 4, utility metrics increase while diversity metrics decrease with larger layers. This trade-off aligns with the model’s design: larger layers focus on utility and specific interests for exploitation, while smaller layers prioritize diversity for exploration. This structure provides users with a flexible mechanism to adjust recommendations from niche, relevant items to broader, exploratory sets based on their preferences. A case study is provided in Appendix C.3.
To further validate the hierarchical structure’s capability to capture varying degrees of group preferences, we analyzed the average norm of node embeddings across layers (Figure 3). The results reveal a consistent increase at larger layers (farther from the center in hyperbolic space). As previously mentioned, hyperbolic geometry ensures that popular items or exploratory users are positioned closer to the origin, while niche items or users are positioned toward the boundaries. This observation aligns with the theoretical properties of hyperbolic space and supports our hierarchical design: higher-level (smaller-layer) nodes effectively encapsulate abstract and aggregated preferences, whereas lower-level (larger-layer) nodes focus on fine-grained, personalized preferences.
6 CONCLUSION
In this work, we present HERec, a hyperbolic graph-llm framework that effectively balances exploration and exploitation in recommender systems. Our approach integrates semantic and collaborative information within hyperbolic space, leveraging the unique properties of hyperbolic geometry, supported by rigorous mathematical proof. Furthermore, we introduce a personalized mechanism for balance the exploration-exploitation trade-off using a hierarchical representation structure, empowering users to adjust recommendation preferences flexibly. Extensive experiments demonstrate that HERec achieves state-of-the-art performance, delivering up to a $5.49\%$ improvement in utility metrics and an $11.39\%$ increase in diversity metrics. Remarkably, HERec is the first framework to excel in both utility and diversity simultaneously, marking a significant advancement in the field.
Appendix A Proof
In this section, we provide a rigorous mathematical proof of Proposition 1, demonstrating the unique advantage of the aligning mechanism in the hyperbolic space.
A.1 Gradient Analysis in Hyperbolic Space.
Suppose $\mathbf{x}$ is a hyperbolic embedding derived from  4.1, and $\mathbf{y}$ is the corresponding semantic embedding from  4.2, the hyperbolic distance is given by:
$d_{\mathcal{H}}(\mathbf{x},\mathbf{y})=\operatorname{arccosh}(z)$ (15)
where $z=-\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}$.
For large $\|\mathbf{x}\|$ and $\|\mathbf{y}\|$, given $x_{0}=\sqrt{1+\|\mathbf{x}\|^{2}}$ and $y_{0}=\sqrt{1+\|\mathbf{y}\|^{2}}$, we adopt the following approximations for further calculation:
$x_{0}\approx\|\mathbf{x}\|,\quad y_{0}\approx\|\mathbf{y}\|,\quad\sqrt{z^{2}-1}\approx z$ (16)
Thus, the Lorentz product can be approximated as:
$z=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}\approx\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)$ (17)
Let $x_{j}=\|\mathbf{x}\|\hat{x}_{j}$ and $y_{j}=\|\mathbf{y}\|\hat{y}_{j}$ (where $\hat{x}_{j}$ represents the direction of $x_{j}$). The partial derivative of $z$ with respect to $x_{j}$ is calculated as follows:
$\frac{\partial z}{\partial x_{j}}$ $=\frac{x_{j}}{x_{0}}y_{0}-y_{j}$ (18)
$\approx\frac{x_{j}}{\|\mathbf{x}\|}\|\mathbf{y}\|-y_{j}$
$=\|\mathbf{y}\|(\hat{x}_{j}-\hat{y}_{j})$
Therefore, the gradient norm of $d_{\mathcal{H}}$ for $\mathbf{x}$ can be approximated by:
$\|\nabla_{\mathbf{x}}d_{\mathcal{H}}\|$ $=\frac{\|\nabla_{\mathbf{x}}z\|}{\sqrt{z^{2}-1}}$ (19)
$\approx\frac{\|\nabla_{\mathbf{x}}z\|}{z}$
$=\frac{\|\mathbf{y}\|\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)}$
$=\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}$
This result suggests that nodes with large norms experience smaller gradient updates, preserving local structures, while nodes with smaller norms undergo larger updates, refining global hierarchical relationships and effectively capturing hierarchical structures. This aligns with the properties of hyperbolic geometry, where small positional changes near the origin result in significant distance changes, and vice versa. Our dynamic gradient update mechanism is well-suited to this characteristic, enabling precise adjustments and enhancing the accuracy of the updates.
A.2 Gradient Analysis Euclidean Space.
For comparison, we examine the gradient behavior in Euclidean space. The Euclidean distance is defined as:
$f(\mathbf{x})=d_{E}(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|$ (20)
The gradient of $f(\mathbf{x})$ with respect to $\mathbf{x}$ is:
$\nabla_{\mathbf{x}}f(\mathbf{x})=\frac{\mathbf{x}-\mathbf{y}}{\|\mathbf{x}-\mathbf{y}\|}=\frac{\mathbf{x}-\mathbf{y}}{d_{E}(\mathbf{x},\mathbf{y})}$ (21)
which has a magnitude of:
$\|\nabla_{\mathbf{x}}f(\mathbf{x})\|=1$ (22)
The constant gradient magnitude demonstrates Euclidean space’s uniform update behavior, lacking intrinsic adaptation to hierarchical structures. This fundamental difference in gradient dynamics explains hyperbolic space’s superiority for hierarchical representation learning.
A.3 Error Bound Analysis
The approximations used in the proof rely on the assumption that $|x|$ and $|y|$ are large, which is common in high-dimensional hyperbolic embedding spaces. In practice, with embedding dimensions such as 50 (as used throughout our experiments), the norms tend to be sufficiently large for the approximation to hold effectively. To further substantiate this assumption, we calculated the error bound as follows:
The exact Lorentz inner product is given by:
$z=\sqrt{1+|x|^{2}}\sqrt{1+|y|^{2}}-|x||y|\cos\theta,$
and is approximated as:
$z_{\text{approx}}=|x||y|(1-\cos\theta).$
Using a Taylor expansion, we have:
$\sqrt{1+|x|^{2}}=|x|\left(1+\frac{1}{2|x|^{2}}-\frac{1}{8|x|^{4}}+\cdots\right).$
Thus, the absolute error becomes:
$|z-z_{\text{approx}}|=\frac{|y|}{2|x|}+\frac{|x|}{2|y|}.$
The relative error in the Lorentz product is bounded by:
$\epsilon_{z}\leq\frac{1}{2}\left(\frac{1}{|x|^{2}}+\frac{1}{|y|^{2}}\right)\frac{1}{1-\cos\theta}.$
Additionally, the approximation error in $\sqrt{z^{2}-1}\approx z$ is given by:
$\sqrt{z^{2}-1}=z\left(1-\frac{1}{2z^{2}}+\cdots\right),$
which implies:
Relative $\text{Relative error}\leq\frac{1}{2z^{2}}.$
Combining these results, the overall error bound is:
Relative Error in |$\text{Relative Error in }|\nabla d_{H}(x,y)|\leq\frac{1}{2|x|^{2}(1-\cos\theta)}+\frac{1}{2z^{2}}.$
For instance, with $|x|=|y|=5$ and $\theta=\frac{\pi}{2}$, the total error is approximately $2.1\%$, demonstrating high approximation accuracy in our setting.
Appendix B Experimental Settings
In this section, we present details of our experimental settings, the distribution of head and tail items (Table 3), implementation details, and an explanation of the diversity metrics used. Notably, the high proportion of tail users (T80) in these datasets highlights the prevalence of long-tail distributions, reaffirming the assumption of a power-law distribution commonly observed in real-world data.
B.1 Implementation Details
We use an embedding dimension of 50 and a training batch size of 1024. The embeddings in hyperbolic space are optimized using Riemannian SGD (RSGD) [2, 25], while the MLP adapter is trained with the Adam optimizer [13]. Training employs early stopping based on Recall@10, with a patience of 10 epochs. User and item profiles are generated using GPT-3.5-turbo and encoded with text-embedding-ada-002 [22] for the Amazon-Books and Yelp datasets, and BERT [8] for Google-Reviews. All experiments are conducted on NVIDIA A40/A100 GPUs with 40GB of memory.
For the final experiments, we perform a grid search over commonly used hyperparameters: learning rate $\in\{1\mathrm{e}{-3},5\mathrm{e}{-3},1\mathrm{e}{-2},5\mathrm{e}{-2},1\mathrm{e}{-1}\}$, weight decay $\in\{1\mathrm{e}{-3},1\mathrm{e}{-2},1\mathrm{e}{-1}\}$, number of negative samples $\in\{10,20,50\}$, and alignment weight $\in\{1\mathrm{e}{-3},1\mathrm{e}{-2},1\mathrm{e}{-1}\}$. A full list of tuned hyperparameters is available in our attached GitHub repository.
Table 3: Statistics of the experimental data.
Dataset #User #Item #Interactions Density
All H20(%) T80(%)
[…]
B.2 Diversity Metrics
Distance Diversity (Div) quantifies the variation among recommended items, encouraging diversity within the recommendation:
$Div(u)=\frac{\sum_{i\in R_{u}}\sum_{j\in R_{u},j\neq i}d(i,j)}{\text{Number of}(i,j)\text{ pairs}}$ (23)
where $R_{u}$ is the recommendation list for user $u$, and $d(i,j)$ represents the distance between items $i$ and $j$. Number of $(i,j)$ pairs is calculated as $\binom{|R_{u}|}{2}=\frac{|R_{u}|(|R_{u}|-1)}{2}$. For fair comparison, we compute distances within the same representation space across models.
Shannon Entropy (H) quantifies the unpredictability within the recommendations, with higher entropy signifying greater diversity:
$H=-\sum_{i\in\text{set}(\sum_{u}R_{u})}p(i)\log p(i)$ (24)
where $\sum_{u}R_{u}$ denotes all recommended items across users (including duplicates), and $p(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}R_{u})}\text{count}(i)$ is the proportion of item $i$ among all recommended items.
Expected Popularity Complement (EPC) assesses the extent of bias towards popular items by promoting niche recommendations:
$EPC=1-\frac{1}{|\text{set}(\sum R_{u})|}\sum_{i\in\text{set}(\sum_{u}R_{u})}\frac{\text{pop}(i)}{\max_{i}(\text{pop}(i))}$ (25)
where $N_{u}$ is the set of items user $u$ has interacted with, and $\text{pop}(i)=\text{count}(i)/\sum_{i\in\text{set}(\sum_{u}N_{u})}\text{count}(i)$ denotes item $i$’s popularity.
Appendix C Supplement to Hyperbolic Hierarchical Clustering
In this section, we provide details of the hyperbolic hierarchical clustering algorithm, together with complexity analysis and a detailed case study.
C.1 Hyperbolic Hierarchical Clustering
Algorithm 1 describes the hyperbolic hierarchical clustering procedure used to organize hyperbolic embeddings into a multi-layer structure. The algorithm takes as input the hyperbolic embeddings $X$ and a layer proportion parameter $k$, which controls the reduction rate of points per hierarchical layer.
The maximum number of layers $L$ is set to $\log_{k}(|X|)$, ensuring that each subsequent layer contains approximately $1/k$ times the points of the previous layer. Starting from the bottom layer containing all points, the algorithm iteratively constructs higher layers in a bottom-up fashion.
At each layer $l$, cluster centroids are randomly initialized in hyperbolic space, with the number of clusters equal to $|D_{l}|/k$, where $|D_{l}|$ is the number of points in the current layer. Clustering proceeds by repeatedly assigning each point to its nearest centroid (in terms of hyperbolic distance) and updating each centroid to the hyperbolic mean of its assigned points. This process continues until the centroids converge.
Upon convergence, the centroids from the current layer serve as the points for the next upper layer. By repeating this process, the algorithm constructs a hierarchy of clusters, effectively capturing both the geometric structure and hierarchical relationships inherent in hyperbolic embeddings.
Algorithm 1 Hyperbolic Hierarchical Clustering
  Input: Hyperbolic embeddings $X$, layer proportion $k$
  Set the maximum number of layers $L=\log_{k}(|X|)$
  Initialize the points of bottom layer $D_{L}=X$
  for $l=L$ to $1$ (bottom to top) do
     Randomly initialize cluster centroids in the hyperbolic space $M_{l}=\{\mu_{l}\ |\ \mu_{l}\in\mathbb{H}\}$, $|M_{l}|=|D_{l}|/k$
     repeat
        for each point $d_{l,i}\in D_{l}$ do
           Assign $d_{i}$ to the cluster with the nearest centroid:
           $c_{l,i}=\arg\min_{\mu\in M_{l}}d_{\mathcal{H}}(d_{i},\mu)$
        end for
        for each cluster $c_{l,i}\in C_{l}$ do
           Update the centroid $\mu_{l,i}$ as the hyperbolic mean $m_{\mathcal{H}}$ of all points assigned to cluster $c_{l,i}$:
           $\mu_{l,i}=m_{\mathcal{H}}(\{d\ |\ d\in C_{l,i}\})$
        end for
     until Centroids $M_{l}$ do not change (convergence)
     Initialize the points of the upper layer $D_{l-1}=M_{l}$
  end for
C.2 Complexity Analysis
In this subsection, we analyze the complexity of constructing and applying the hierarchical representation structure. Let $N$ be the total number of users and items, and $k$ be the number of recommendations per user.
Hierarchical Clustering. Building the hierarchy tree involves iterative k-means clustering in hyperbolic space. To optimize efficiency, we assign cluster nodes only to real node embeddings. First, we compute and store pairwise distances, requiring $O(N^{2}/2)$. Each k-means iteration takes $O(i)$, and with $O(\log N)$ levels, the overall complexity of constructing the hierarchy tree is $O(N^{2}/2+i\log N)$.
Recommendation Retrieval & Replacement. Finding the top-$k$ recommendations requires $O(k\log N)$ using nearest neighbour search in hyperbolic space. Adjusting recommendations by retrieving ancestor clusters at layer $l$ requires $O(k)$ time, as it involves only tree traversal.
The proposed hierarchical framework achieves efficient complexity scaling while maintaining flexibility in recommendation adjustments. The hierarchical clustering process, optimized for hyperbolic space, ensures a computationally feasible structure with $O(N^{2}/2+i\log N)$ complexity. The recommendation retrieval and replacement steps remain lightweight, running in $O(k\log N)$ and $O(k)$, respectively. This design enables HERec to dynamically balance exploration and exploitation while ensuring scalability for large-scale recommender systems.
C.3 Hierarchical Structure Case Study
In this subsection, we present a case study examining the exploration-exploitation trade-off in the hyperbolic hierarchical structure. We randomly select a user and identify items that share different layers of the lowest common ancestor (LCA) with the user in the hierarchy tree. Ideally, a larger LCA layer indicates that the user and item were grouped earlier in the hierarchy, suggesting a stronger likelihood of interaction. Conversely, a smaller LCA layer implies that the user and item were grouped only after multiple iterations, indicating a weaker likelihood of interaction.
 User Profile:
The user enjoys books with supernatural elements, horror and paranormal romance. They also appreciate fast-paced action and suspenseful thrillers with political intrigue and espionage. The user likes complex characters dealing with dark themes.
Item 1 Profile (LCA=10):
The book ’Spirit Fighter (Son of Angels, Jonah Stone)’ would likely appeal to young readers who enjoy stories about supernatural beings and the fight between good and evil. It is a book that offers entertainment and inspiration, as well as a glimpse into the supernatural world.
Item 2 Profile (LCA=5):
Fans of contemporary romance novels will enjoy Barefoot Summer (A Chapel Springs Romance). This book combines heartwarming romance, relatable heartaches, and lighthearted humor to create a delightful and engaging story. It is ideal for those looking for an easy-to-read and entertaining summer novel.
As demonstrated in the case study, the user shows a preference for supernatural elements, which aligns closely with the description of item 1, indicating a high likelihood of interaction. In contrast, while the user profile mentions a preference for dark themes, item 2 focuses on romance and heartwarming narratives, suggesting a low probability of user interest in item 2.
This analysis aligns seamlessly with the intended design of our hierarchical tree, further validating the effectiveness of our model in capturing and leveraging nuanced user-item relationships.
Table 4: Comparison with feature-enhanced Euclidean baselines, with the best-performing results highlighted in bold.
Data Amazon-books Yelp Google-reviews
Metrics R@20 N@20 Div@20 H@20 EPC@20 R@20 N@20 Div@20 H@20 EPC@20 R@20 N@20 Div@20 H@20 EPC@20
LightGCN+ 0.1466 0.0901 0.4349 10.3915 0.7802 0.1156 0.0733 0.2697 9.6903 0.7526 0.1438 0.0940 0.2707 10.8664 0.7801
LightGCL+ 0.1545 0.0992 0.5064 11.9761 0.8728 0.1131 0.0739 0.2759 11.8617 0.8581 0.1443 0.0946 0.2744 12.4700 0.8776
SimGCL+ 0.1494 0.0926 0.5163 11.4751 0.8342 0.1195 0.0752 0.3152 9.9952 0.7811 0.1504 0.0982 0.3030 11.2111 0.7953
HCCF+ 0.1525 0.0940 0.4316 11.7984 0.8616 0.1170 0.0729 0.2884 11.6906 0.8577 0.1438 0.0941 0.2524 11.5433 0.8406
Ours 0.1682 0.1069 0.5754 12.2543 0.8758 0.1224 0.0762 0.3961 12.3290 0.8773 0.1598 0.0992 0.3543 12.7690 0.8728
Table 5: Standard deviation of key metrics across datasets over 10 independent runs.
Dataset Recall@20 NDCG@20 DIV@20 H@20 EPC@20
Amazon 0.0010 0.0004 0.0067 0.0035 0.0002
Yelp 0.0009 0.0004 0.0091 0.0017 0.0014
Google 0.0010 0.0003 0.0077 0.0008 0.0009
Table 6: Direct message passing within hyperbolic space.
Dataset Recall@20 NDCG@20 DIV@20 H@20 EPC@20
Amazon 0.1721 0.1082 0.5760 12.211 0.8773
Yelp 0.1205 0.0793 0.4005 12.321 0.8755
Google 0.1611 0.0988 0.3532 12.766 0.8731
Appendix D More Experimental Results
In this section, we will show more experimental results to further analyze the flexibility and effectiveness of our HERec.
D.1 Baselines with Feature Enhancement
For further validation, we assessed the impact of feature integration on Euclidean baselines. By incorporating semantic embeddings into the Euclidean space—mirroring our approach in Section 4.2—we created enhanced variants (denoted by “+”). As reported in Table 4, these models generally achieve higher utility scores but do not consistently improve diversity. This reinforces our main findings: while feature integration benefits utility in Euclidean frameworks, it falls short for diversity, highlighting the unique strengths of hyperbolic alignment detailed in Section 4.4.
D.2 Statistical Significance Testing
To ensure the robustness and reliability of our results, we conducted 10 independent runs for each experimental setting and reported the standard deviation for all key metrics. Results are shown in Table 5. Across all datasets and evaluation metrics, we observed consistently low standard deviations, indicating high stability and reproducibility of our findings. This statistical rigor further substantiates the significance of our experimental conclusions.
D.3 Alternative Hyperbolic Message Passing
We also explored an alternative message passing strategy by replacing the original hyperbolic graph collaborative filtering mechanism with direct message passing in hyperbolic space, thereby eliminating the need for tangent space projections. This variant achieves performance comparable to our primary method, results can be found in Table 6. These findings suggest that direct message passing in hyperbolic space is not only a viable alternative but could also serve as a complementary component within our framework, potentially offering additional flexibility for future extensions.
Appendix E Profile Decoder
System Instruction: Describe what types of items this user is likely to enjoy./ Describe what types of users would enjoy this item. Input Prompt: <USER_EMBED>/<ITEM_EMBED> Output: The user would prefer items which …/ The item would be preferred by users who …
Figure 5: A depiction of model prompt instruction.
To further evaluate the effectiveness of hyperbolic alignment and the expressiveness of our learned hyperbolic embeddings, we introduce a profile decoder designed to reconstruct the original profile from these embeddings, thereby broadening their application beyond recommendation tasks. Our method begins by mapping the hyperbolic embeddings back into Euclidean space to facilitate alignment with traditional network layers. These Euclidean-mapped embeddings are then processed through a mixture of experts (MoE) module. The MoE serves to refine the embeddings, ensuring consistency in both dimensionality and semantic representation [20]. Once processed, these embeddings are fed into the language model, specifically after the positional layer, to incorporate them within a sequential architecture. To enhance training efficiency and minimize computational cost, we freeze the language model’s parameters, making the MoE module the only trainable component in the architecture.
[…]
E.1 Decoder Quality
To assess the effectiveness of our decoder mechanism, we compare the embeddings from our model against those from Euclidean embeddings (LightGCN) and Hyperbolic embeddings (HICF) as inputs to the decoder. The results are shown in Table 7. We use BERTScore [44] to evaluate the semantic alignment of generated sentences with the original profile. Results indicate that our HERec outperforms both Euclidean and hyperbolic baselines, demonstrating that our model’s embeddings support recommendation and retain user/item profile information. This finding underscores that our model captures collaborative information alongside recoverable language semantics: a capability not achieved by previous models.
E.2 Decoder Case Study
To vividly illustrate the semantic capacity of our embeddings, we present two case studies demonstrating the model’s ability to recover user and item profiles.
Original user profile:
Based on the user’s interactions and reviews, this user is likely to enjoy unique dining experiences with a variety of options, trendy atmosphere with good drink and food selections, premium quality frozen yogurt with a wide variety of toppings, creative burgers and hot dogs, delicious Thai cuisine, and authentic and spicy Chinese cuisine served family-style.
Recovered user profile:
Based on the user’s interactions, it is likely that they would enjoy businesses that offer a variety of cuisines, including Mexican, Italian, and Asian fusion. They appreciate good food, drinks, and a relaxed atmosphere. They also value quality and freshness in their food.
In this example, the profile recovered by our decoder accurately captures the user’s preferences for Thai and Chinese cuisines, broadly summarized as an affinity for Asian fusion flavours. Additionally, our decoder identifies the user’s interest in diverse culinary experiences, indicating a propensity toward exploration.
Original item profile: Ceol Irish Pub in Reno is a popular bar and pub known for its happy hour deals, great service, and friendly atmosphere. It is suitable for those who enjoy trivia nights, live music, and a wide selection of drinks, particularly whiskey. The pub has a welcoming environment, with bartenders who are attentive and quick to serve.
Recovered item profile:
Users who enjoy a variety of beer options, live music, and a relaxed atmosphere would enjoy Ceol Irish Pub.
The alignment between the original and recovered item profiles shows that our decoder effectively captures key thematic elements of Ceol Irish Pub. The original profile emphasizes aspects such as happy hour, a wide selection of drinks, live music, and a welcoming atmosphere with attentive service. The recovered profile distills these characteristics, focusing on the pub’s appeal for users who enjoy diverse beer options, live music, and a relaxed setting. This alignment suggests that our decoder reliably summarizes key attributes, albeit with a degree of abstraction.
Table 7: Experiments assessing decoder quality.
Data Amazon-books Yelp
BERTScore P R F1 P R F1
Euclidean 0.3266 0.2854 0.3064 0.3511 0.2238 0.2872
Hyperbolic 0.3498 0.3009 0.3257 0.3608 0.2398 0.3021
Ours 0.3665 0.3167 0.3418 0.3716 0.2453 0.3081